### v1

# Swarming in the Wild: A Distributed Communication-less Lloyd-based Algorithm dealing with Uncertainties

Abstract
In this work, we present a distributed algorithm for swarming in complex environments that operates with no communication, no a priori information about the environment, and using only onboard sensing and computation capabilities. We provide sufficient conditions to guarantee that each robot reaches its goal region in a finite time, avoiding collisions with obstacles and other robots without exceeding a desired maximum distance from a predefined set of neighbors (flocking constraint). In addition, we show how the proposed algorithm can deal with tracking errors and onboard sensing errors without violating safety and proximity constraints, still providing the conditions for having convergence towards the goal region. To validate the approach, we provide experiments in the field. We tested our algorithm in GNSS-denied environments i.e., a dense forest, where fully autonomous aerial robots swarmed safely to the desired destinations, by relying only on onboard sensors, i.e., without a communication network. This work marks the initial deployment of a fully distributed system where there is no communication between the robots, nor reliance on any global localization system, which at the same time it ensures safety and convergence towards the goal within such complex environments.
Index Terms:
Multi-robot systems, Distributed control, Lloyd-based algorithms.
Video—https://mrs.fel.cvut.cz/rbl-wild-2024
Github—https://github.com/ctu-mrs/rbl_controller
I Introduction
Over the past few decades, numerous applications have emerged for multi-robot systems, exhibiting their undisputed benefits in diverse fields. These applications span domains such as logistics, surveillance, environment monitoring, disaster response, search and rescue, and agriculture among others, demonstrating their versatility. For many of these applications, having a swarm of robots able to flock together towards a desired location can be highly beneficial, due to redundancies and parallelization. The definition of flocking is provided by the seminal work of Reynolds [1], which defines it as a non-colliding aggregate motion. Depending on the environments where robots operate, specific factors like the reliability or availability of direct communication between robots may be compromised.
In this work, we propose a Lloyd-based algorithm that relies only on on-board sensors. We aim to synthesize a distributed control algorithm, where each robot safely reaches the desired position, maintaining proximity between robots in the team. In particular, we provide sufficient conditions to guarantee safety, proximity maintenance and convergence towards the goal in cluttered environments by accounting for bounded tracking errors and uncertainties in the estimation of neighbors and obstacles positions.
Our algorithm takes inspiration from nature; in fact, just like flock of birds or a school of fish, the robots do not need a leader and each robot makes its own decisions based on relative information, without relying on explicit communication. We validate our approach through an extensive simulation and experimental campaign in the field. In Figure 1, we illustrate a snapshot during a mission in the forest. This work represents the first deployment of a fully distributed system with absence of any communication between the robots and any global localization system, which also provides conditions for safety and convergence towards the goal in such a complex environments.
Figure 1: Deployment of the proposed algorithm in a real-world scenario with $3$ UAVs navigating in a dense forest.
I-A Related works
Prior literature provides a considerable amount of algorithms for distributed motion coordination of multiple robots, but only a few of them can manage to work in a cluttered environment by also imposing the flocking constraint. Among them, a smaller subset is devoted to applications in the field without relying on direct communication between the robots.
To position our algorithm within the state of the art, we considered the following features:
  • 1.
    Distributed control: each robot makes decisions on the basis of local information, which is obtained only by relying on onboard sensors;
  • 2.
    Asynchronous control: each robot makes decisions independently, in asynchronous fashion;
  • 3.
    Communication independence: the robots do not need to communicate, since the algorithm does not rely on unobservable information, e.g., future trajectories, velocities, and accelerations;
  • 4.
    Safety: the system is provably safe, i.e., given the upper bounds on uncertainties (see item 7), collision and obstacle avoidance is guaranteed at any time;
  • 5.
    Convergence: the sufficient conditions for the robots to converge to their goal locations are provided (i.e., absence of deadlock and live-lock);
  • 6.
    Proximity maintenance: the algorithm allows to set hard constraints on the maximum distance between robots;
  • 7.
    Uncertainties: the system is provably safe and respects proximity constraints even in the presence of bounded tracking errors and/or bounded measurement uncertainties;
  • 8.
    Cluttered and highly dense environments: the algorithm can safely manage the presence of static obstacles in the environment. It can also handle highly dense cluttered environments, such as a dense forest.
  • 9.
    Onboard localization: the algorithm is tested by relying on onboard localization in unknown environments.
  • 10.
    Experiments: the proposed algorithm is tested on real robotic platforms in unstructured environments;
  • 11.
    Open source code: the algorithm is open source on Github.
Method [2] [3] [4] [5] [6] [7] [8] [9] [10] [11] [12] [13] [14] [15] [16] [17] [18] [19] [20] [21] [22] [23] [24] [25] [26] ours
Distributed $\circ$ $\circ$ $\circ$ o $\circ$ $\times$
Asynchronous $\circ$ $\circ$ $\circ$ $\times$ $\times$ $\times$ $\times$ $\times$
w/o Comm. $\circ$ $\times$ $\times$ $\times$ $\times$ $\times$ $\circ$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Safety $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\times$
Convergence $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Uncertainties $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\times$ $\circ$ $\circ$ $\circ$
Prox. maint. $\times$ $\times$ $\circ$ $\circ$ $\circ$ $\circ$ $\times$ $\times$ $\times$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\times$ $\times$ $\times$ $\circ$ $\times$
Clutt. Envir. $\times$ $\circ$ $\circ$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
On-board Loc. $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Experiments $\times$ $\circ$ $\circ$ $\circ$ $\times$
Open Source Code $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Table I: List of state of the art algorithms. We indicate with “$\times$” if the algorithm does not implement the feature, “✓” if it does, and “o” if it does only partially.
In Table I, we report an overview of the most closely related algorithms in the literature. We mark with “$\times$” if the algorithm does not implement the feature, “✓” if it does, and finally, “o” if it does but only partially. As shown in Table I, our algorithm is unmatched in fulfilling all the desired features of interest. In fact, despite [5, 4] proposing valid solutions to the problem, their main limitations are the requirement of a trajectory broadcasting network, the absence of both convergence analysis and a rigorous account of tracking errors and uncertainties. Similarly, in [6] the robots need to broadcast information such as poses, velocities, poses covariance, and global extrinsic transformations. On the other hand, [7, 8], do not rely on a communication network, but they exhibit degraded performance in crowded scenarios as well as they lack of safety, proximity maintenance, convergence guarantees, and robustness to tracking errors and uncertainties. In [10], the authors present a distributed solution, which in theory does not rely on communication and accounts for uncertainties. However, in the experiments they rely on motion capture measurements adding limited noise errors, which is reasonable only in structured environment (i.e., laboratory conditions). In Table  I, we report other state-of-the-art algorithms. Notably, all of them share limitations related to the use of a communication network and/or to not rigorously account for safety, convergence, proximity maintenance and the uncertainties. Conversely in our work, we provide a method which implements all the fore-mentioned features as well as a clear strategy for the selection of the relevant parameters, which in the state-of-the-art is often an unclear or empirical process.
I-B Contribution and paper organization
Our contribution is fivefold:
  • We propose a novel algorithm for flocking control in dense cluttered environments exploiting some of the theoretical findings in [2]. With respect to [3], where the authors introduced the flocking behavior in the Lloyd-based framework, we do not impose the line-of-sight constraint, i.e., two robots are considered in the proximity if they are close enough. This allows to increase the robots mobility and also to not rely on communication. At the same time, it introduces high uncertainties in the position estimation of the neighbors when robots are not in the line of sight.
  • We provide sufficient conditions for safety, proximity maintenance, and convergence even in cluttered environments and with flocking constraints.
  • We propose a novel adaptive method to account for tracking errors and measurement uncertainties.
  • We tailor the use of Convex Weighted Voronoi Diagrams (CWVD) [13] for our problem. By doing that, we considerably improve the performance of the algorithm in dense cluttered environments such as a dense forest.
  • We validate the algorithm in the forest, relying only on on-board sensors i.e., IMU, 2-D Lidar and UVDAR system, without relying neither on a communication network nor GNSS signal.
The paper is organized as follows: Sec. II provides the problem description, in Sec. III we present the background material and the proposed approach, in Sec. IV we provide simulation results and a comparison with a state-of-the-art approach. In Sec. V we validate the proposed approach through experimental results in the field, and finally in Sec. VI we draw the conclusions and the future research directions.
II Problem description
Let us denote the positions of the robots with $p=[p_{1},\dots,p_{N}]^{\top}$, where $p_{i}=[x_{i},y_{i}]^{\top}$ and the obstacle positions by $o=[o_{1},\dots,o_{N_{o}}]^{\top}$, $o_{k}=[x_{k},y_{k}]^{\top}$. The mission space is represented by $\mathcal{Q}$, the weighting function that assess the weights of the points $q\in\mathcal{Q}$ is $\varphi_{i}(q):\mathcal{Q}\rightarrow\mathbb{R}_{+}$, while $\delta_{i}$ and $\delta_{k}^{o}$ are the radius of encumbrance of the $i$–th robot and the $k$–th obstacle respectively.
The problem that we aim to solve reads as follow:
Problem 1.
Let $N$ be the number of robots in the swarm, let $\mathcal{N}_{i}$ be the set of neighbors of the robot $i$, and $\bar{\mathcal{N}}_{i}\subseteq\mathcal{N}_{i}$ the set of neighbors that has to remain connected the entire mission. We want to synthesize a distributed control law, such that the swarm maintains the desired proximity constraints, which is characterized by a set of maximum distances $\Gamma_{ij},\,\,\,\forall i,\forall j\in\bar{\mathcal{N}}_{i}$. Each robot in the swarm has to reach its goal region avoiding collisions with other robots and obstacles.
Assumption 1.
Each robot $i$ can perceive the relative position of robot (obstacle) $j$ if $j\in\mathcal{N}_{i}$ ($j\in\mathcal{O}_{i}$). Where $j\in\mathcal{N}_{i}$ if $\|p_{i}-p_{j}\|\leq R_{s}$ ($j\in\mathcal{O}_{i}$ if $\|p_{i}-o_{j}\|\leq R_{s}$), with $R_{s}$ indicating the sensing range, equal for all the robots.
Assumption 2.
The flocking constraints are symmetric, i.e., $\|p_{i}-p_{j}\|\leq\Gamma_{ij},\forall j\in\bar{\mathcal{N}}_{i},\forall i=1,\dots,N$, where $\Gamma_{ij}=\Gamma_{ji}<R_{s}$.
Assumption 3.
The size of the neighboring robots and obstacles is known and it can be approximated with circles/cylinders.
Assumption 4.
In our theoretical analysis, we assume that the robots are holonomic agents.
Assumption 5.
From Section III-C, we assume more complex dynamical systems and a more realistic perception system, where the tracking error is bounded as well as the error on the measured distances from neighbors and obstacles.
III Proposed Approach
The presented approach is grounded in the Lloyd algorithm [27]. In the literature, its distributed nature was exploited to develop novel algorithms for several collective behaviors [3, 2, 13, 28, 29].
Let us consider the following cost function
[…]
where $\mathcal{V}_{i}=\{\mathcal{V}_{i}^{p}\cap\mathcal{V}_{i}^{o}\}$ with $\mathcal{V}_{i}^{p}$ being the $i$–th Convex Weighted Voronoi Diagram (CWVD) introduced in [13], which accounts only for neighboring robots positions
$\small\begin{split}\mathcal{V}_{i}^{p}=\{q\in\mathcal{Q}\mid w_{i}^{x}\cos\alpha_{ij}+w_{i}^{y}\sin\alpha_{ij}<\frac{1}{\epsilon_{p}}&\left\|p_{i}-p_{j}\right\|,\\ &\forall j\in\mathcal{N}_{i}\}.\end{split}$ (2)
While $\mathcal{V}_{i}^{o}$ is the $i$–th CWVD accounting only for neighboring obstacles positions
$\small\begin{split}\mathcal{V}_{i}^{o}=\{q\in\mathcal{Q}\mid w_{i}^{x}\cos\alpha_{ij}+w_{i}^{y}\sin\alpha_{ij}<\frac{1}{\epsilon_{o}}&\left\|p_{i}-o_{j}\right\|,\\ &\forall j\in\mathcal{O}_{i}\},\end{split}$ (3)
where $w=[w^{x},w^{y}]^{\top}=q-p_{i}$ and $\alpha_{ij}=\arctan\left(\frac{y_{j}-y_{i}}{x_{j}-x_{i}}\right)$, while $\epsilon_{p(o)}\in[1,2]$, is an arbitrary real number, which regulates the cautiousness in the robot motion (i.e., $\epsilon_{p(o)}=1$ more aggressive behavior, $\epsilon_{p(o)}=2$ more cautelative behavior, notice that the latter coincides with the classic Voronoi diagram).
Figure 2: Representation of the Convex Weighted Voronoi Diagram (CWVD) $\mathcal{V}_{i}^{p}$ in (2) with $\epsilon_{p}=2$ and $\epsilon_{p}=1$. Where the green circle is the i–th robot and the cyan circles represents the neighboring robots.
In Figure 2, we depict an example of the CWVD (2) for the case where $\epsilon_{p}=2$ and $\epsilon_{p}=1$.
By exploiting Assumption 4, we can impose $\dot{p}_{i}=u_{i}$. By following the gradient descent of the cost function (1), we obtain the following proportional control law
$\dot{p}_{i}\left(\mathcal{V}_{i}\right)=-k_{p,i}\left(p_{i}-c_{\mathcal{V}_{i}}\right),$ (4)
where $k_{p,i}>0$ is a positive real number, and
[…]
is the centroid position evaluated over the $\mathcal{V}_{i}$ cell. By imposing (4), assuming a time-invariant $\varphi_{i}(q)$, each robot $i$, $\forall i,1\dots,N$, converges asymptotically to the centroid position $c_{\mathcal{V}_{i}}$ [30]. In the following, we show how, by introducing suitable modifications to the geometry of cell $\mathcal{V}_{i}$ and crafting an appropriate weighting function $\varphi_{i}(q)$, we can attain a safe and effective algorithm for multi-robot swarming.
III-A Cell geometry $\mathcal{F}_{i}$
The reshape of the cell geometry is both necessary and sufficient to ensure safety guarantees. Hence, we define the cell $\mathcal{A}_{i}=\{\tilde{\mathcal{V}}_{i}^{p}\cap\tilde{\mathcal{V}}_{i}^{o}\cap\mathcal{S}_{i}\}$, where
$\tilde{\mathcal{V}}_{i}^{p}=\left\{\begin{array}[]{l}\left\{q\in\mathcal{V}_{i}^{p}\right\}\text{ if }\frac{1}{2}\left\|p_{i}-p_{j}\right\|\geq\Delta_{ij}\\ \left\{q\in\mathcal{Q}\mid\left\|q-p_{i}\right\|\leq\left\|q-\tilde{p}_{j}\right\|\right\}\text{ otherwise, }\end{array}\right.$ (6)
with $j\in\mathcal{N}_{i}$, $\Delta_{ij}=\delta_{j}+\delta_{i}$, and $\tilde{p}_{j}=p_{j}+2\left(\Delta_{ij}-\frac{\left\|p_{i}-p_{j}\right\|}{2}\right)\frac{p_{i}-p_{j}}{\left\|p_{i}-p_{j}\right\|}$, takes into consideration the physical dimension of the robots. On the other hand,
$\tilde{\mathcal{V}}_{i}^{o}=\left\{\begin{array}[]{l}\left\{q\in\mathcal{V}_{i}^{o}\right\}\text{ if }\frac{1}{2}\left\|p_{i}-p_{j}\right\|\geq\Delta^{o}_{ij}\\ \left\{q\in\mathcal{Q}\mid\left\|q-p_{i}\right\|\leq\left\|q-\tilde{o}_{j}\right\|\right\}\text{ otherwise, }\end{array}\right.$ (7)
with $j\in\mathcal{O}_{i}$, $\Delta^{o}_{ij}=\delta_{i}+\delta^{o}_{j}$ and $\tilde{o}_{j}=o_{j}+2\left(\Delta^{o}_{ij}-\frac{\left\|p_{i}-o_{j}\right\|}{2}\right)\frac{p_{i}-o_{j}}{\left\|p_{i}-o_{j}\right\|}$, takes into consideration the size of the robot and the obstacles, while
$\mathcal{S}_{i}=\left\{q\in\mathcal{Q}\mid\left\|q-p_{i}\right\|\leq r_{s,i}\right\},$ (8)
with $r_{s,i}$ is half of the sensing radius $R_{s}$.
Theorem 1 (Safety [2]).
By imposing the control $\dot{p}_{i}(\mathcal{A}_{i})$ in (4), collision avoidance is guaranteed at every instant of time.
Proof.
The proof can be found in [2], the addition of static obstacles and CVWD does not change the result. In particular, by adding static obstacles and/or introducing CVWD does not alter the convexity of the set. ∎
We further apply an additional reshaping of the cell geometry in order to maintain the proximity, i.e., to obtain swarming. By imposing our final cell geometry $\mathcal{F}_{i}=\{\mathcal{A}_{i}\cap\mathcal{M}_{i}\}$, where
$\mathcal{M}_{i}=\{q\in\mathcal{Q}\mid\|q-p_{j}\|<\Gamma_{ij},\forall j\in\bar{\mathcal{N}}_{i}\},$ (9)
we guarantee proximity maintenance while preserving safety. In Figure 3, we depict an example of cell geometry $\mathcal{F}_{i}$.
Figure 3: Example of $\mathcal{F}_{i}$ and $\mathcal{A}_{i}$ cell geometry with $\epsilon_{p}=\epsilon_{o}=2$. Dashed lines indicate the proximity constraints. Obstacles are represented as grey circles, while the robots are depicted in cyan.
Theorem 2 (Proximity maintenance).
By imposing the control $\dot{p}_{i}(\mathcal{F}_{i})$ in (4), both safety and proximity maintenance are guaranteed at every instant of time.
Proof.
Given that $\mathcal{F}_{i}\subseteq\mathcal{A}_{i}$, safety comes from Theorem 1. Since $\mathcal{F}_{i}$ is convex by definition, then the centroid $c_{\mathcal{F}_{i}}\in\mathcal{F}_{i}$. It implies that by imposing $\dot{p}_{i}(\mathcal{F}_{i})$ in (4), $p_{i}\in\mathcal{F}_{i}$ at any time. Hence, by the definition of the $\mathcal{F}_{i}$ set, proximity maintenance between $i$ and $j$ is always preserved. ∎
III-B The weighting function $\varphi_{i}(q)$
Although ensuring safety and proximity maintenance only involves properly shaping the cell geometry, achieving convergence and enhancing navigation performance largely hinges on how the function $\varphi_{i}(q)$ is defined. We adopt a Laplacian-shaped function centered at $\bar{p}_{i}$, which represents a point in the mission space that ultimately needs to converge to the goal location. We design the $\varphi_{i}(q)$ function as follows:
$\varphi_{i}(q)=\text{exp}{\left(-\frac{\|q-\bar{p}_{i}\|}{\beta_{i}}\right)},$ (10)
where the spreading factor $\beta_{i}\in\mathbb{R}$ is time-variant, it follows the dynamics
$\dot{\beta}_{i}(\mathcal{A}_{i})=\begin{cases}-\beta_{i}\,\,\,&\text{if}\,\,\|c_{\mathcal{A}_{i}}-p_{i}\|<d_{1}\,\land\\ &\|c_{\mathcal{A}_{i}}-c_{\mathcal{S}_{i}}\|>d_{2}\\ -(\beta_{i}-\beta^{D}_{i})&\text{otherwise,}\end{cases}$ (11)
and
$\begin{split}\dot{\bar{p}}_{i}&=\begin{cases}-(\bar{p}_{i}-R(\pi/2-\varepsilon)e_{i})\,\,\,&\text{if}\,\,\|c_{\mathcal{A}_{i}}-p_{i}\|<d_{3}\,\land\\ &\|c_{\mathcal{A}_{i}}-c_{\mathcal{S}_{i}}\|>d_{4}\\ -(\bar{p}_{i}-e_{i})&\text{otherwise,}\end{cases}\\ \bar{p}_{i}&=e_{i}\,\,\,\text{if}\,\,\,\|p_{i}-\bar{c}_{\mathcal{A}_{i}}\|>\|p_{i}-c_{\mathcal{A}_{i}}\|\,\land\,\bar{p}_{i}=R(\pi/2-\varepsilon)e_{i}.\\ \end{split}$ (12)
The parameters $\beta_{i}^{D},d_{1},d_{2},d_{3},d_{4}$ are positive scalars, $R(\theta)$ is the rotation matrix, $e_{i}$ is the goal position for the $i$–th robot, $\varepsilon$ is a small number, which allows to avoid rotations equal to $\pi/2$, while $\bar{{c}}_{\mathcal{A}_{i}}$ is the centroid position computed over the cell $\mathcal{A}_{i}$ with $\bar{p}_{i}\equiv e_{i}$.
Corollary 1 (Convergence in cluttered environment and with flocking constraints).
Let us assume we have an unbounded space cluttered with circular obstacles of radius $\delta^{o}<\Gamma_{ij}/2-2\delta$ and at a minimum distance of $\|o_{i}-o_{j}\|>4\delta+2\delta^{o}$ one to the other. The initial and goal positions are adequately apart from each other, i.e., $\Gamma_{ij}>\|p_{i}(0)-p_{j}(0)\|>\Delta_{ij}$, $\Gamma_{ij}>\|e_{i}-e_{j}\|>\Delta_{ij}$, $\forall i,j=1\dots N$ with $i\neq j$. By imposing the control law $\dot{p}_{i}(\mathcal{F}_{i})$ in (4),  (10), (11), and (12) to compute the centroid position in (5), by properly selecting $d_{1},d_{2},d_{3},d_{4}$, and $\beta^{D}_{i}$, we have that $\lim_{t\rightarrow\infty}\|p_{i}(t)-e_{i}\|<r_{s,i},\forall i=1,\dots,N.$
Figure 4: We depict in cyan the robots and in grey the static obstacles. The crosses represent the final goal position while the dashed lines the proximity constraint topology. In the figure we show two possible deadlock conditions: case 1. is due to the small inter-space between obstacles. case 2. is due to the too short $\Gamma_{ij}$ value (or too big $\delta^{o}$ value).
Proof.
With respect to the proof of convergence with no obstacles and no proximity maintenance constraints, which are provided in [2], we recognize, beside the trivial initial and final conditions requirements, the addition of two potential deadlock scenarios, which are depicted in Figure 4.
Notice that the addition of static obstacles is equivalent to non-cooperative static robots. In order to not affect the proof of convergence obtained in [2], two requirements need to be considered. 1. We require a minimum distance between the obstacles of $\|o_{i}-o_{j}\|>4\delta+2\delta^{o},\,\forall i\neq j$, in order to avoid the presence of narrow hallways that can generate deadlock issues (left hand side of Figure 4). This is a consequence of the results in [2]. In fact, static obstacles can be thought as robots that reached their goal and have a value of $\beta_{i}\rightarrow 0$. However, from [2], by design this is not possible if the robots are at a distance from their goal smaller than $d_{2}$. Since $d_{2}>2\delta$ and $\|e_{i}-e_{j}\|>\Delta_{ij}$, then the condition $\|o_{i}-o_{j}\|>4\delta+2\delta^{o},\,\forall i\neq j$ is equivalent. 2. We require that the maximum radius of the obstacles is $\delta^{o}<\Gamma_{ij}/2-2\delta$, in order to avoid strict proximity maintenance constraints (right hand side of Figure 4). By imposing these additional conditions, the guarantees of convergence in [2] holds. ∎
III-C Dealing with dynamics and uncertainties
The real robot deployment requires to consider a set of additional issues, such as 1. More complex dynamics of the robots must be applied, i.e., $\dot{p}_{i}=u_{i}$ (Assumption 4) cannot be directly applied. 2. Tracking error, which can be due to external disturbances or model mismatches, generating a difference between the desired and the actual velocity, 3. Measurement uncertainties, due to the limitations of the perception system providing an estimate of the positions of other robots and obstacles with a certain error.
To account for the dynamic constraints of the robots, we synthesize the control inputs through an MPC formulation (similarly to [10, 2]), which steers each $i$–th robot towards the centroid position $c_{\mathcal{F}_{i}}$, by constraining the trajectory to not exceed its $\mathcal{F}_{i}$ cell. By doing that, we can account for the robot dynamics without violating the safety and proximity maintenance constraints. Hence, each robot $i$ solves the following optimization problem (for the sake of clarity we discard the $i$ indexing),
$\operatorname*{\min\text{imize}}_{{x}_{k},{u}_{k}}$ $\sum^{N_{t}}_{k=0}J({x}_{k},{u}_{k},c_{\mathcal{F}}),$ (13a)
s.t.: ${x}_{k}=f({x}_{k-1},{u}_{k-1}),\,\,\,\forall k\in(1,N_{t}),$ (13b)
$\,{x}_{k}\in\mathcal{X},\,{u}_{k}\in\mathcal{U},\,\,\,\forall k\in(0,N_{t})$ (13c)
$p_{k}\in\mathcal{F},\,\,\,\forall k\in(0,N_{t})$ (13d)
${x}_{0}={x}_{\textrm{init}},$ (13e)
where $k$ indicates the time index, $N_{t}$ indicates the horizon length, $J$ represents the cost function, ${x}_{k}$ and ${u}_{k}$ are the state and control inputs respectively, $f({x}_{k-1},{u}_{k-1})$ indicates the dynamics of the robot, ${x}_{k}\in\mathcal{X},\,{u}_{k}\in\mathcal{U}$ are the state and input constraints respectively, $p_{k}\in\mathcal{F}$ constrains the future positions of the robot in its cell, and ${x}_{0}={x}_{\textrm{init}}$ is the initial condition.
Both the tracking errors and measurement uncertainties can be accounted for by means of an adaptive strategy. In fact, it can be achieved by adapting the value of $\beta_{i}$ in Equation (11) in an online fashion. An alternative approach would be reshaping the $\mathcal{F}_{i}$ cell, similarly to [10]. However, this method cannot be applied in our case, since, by following this approach, the robot may fall out of its cell due to tracking error, and once it happens, both the cell and the centroid position are no longer defined. Conversely, the value of $\beta_{i}$ defines the aggressiveness of the $i$–th robot. A smaller value corresponds to a centroid position closer to the goal (and hence closer to the boundaries of the $\mathcal{F}_{i}$ cell, namely $\partial\mathcal{F}_{i}$), while larger values of $\beta_{i}$ push the centroid position as far as possible from the boundaries of the cell $\partial\mathcal{F}_{i}$. Thus, managing errors and uncertainties can be achieved by selecting a value of $\beta_{i}$ such that the centroid position stays at a controlled distance from the boundary of the cell. This distance must be chosen in accordance with the maximum positional tracking error and the maximum error in the neighbors’ position estimates. More formally, it can be done by imposing a lower bound to $\beta_{i}$ in (11) equal to
$\beta^{\min}_{i}=\operatorname*{arg\,min}_{\beta_{i}}\left(\|c_{\mathcal{F}_{i}}-p_{\partial\mathcal{F}_{i}}\|-d_{u}\right)^{2},$
where $p_{\partial\mathcal{F}_{i}}$ is the position of the closest point on the contour of $\mathcal{F}_{i}$, while $d_{u}$ is a threshold value that has to be tuned on the basis of the tracking errors and measurement uncertainties. For the sake of clarity, we depicted three scenarios in Figure 5. Figure 5a considers the tracking error; in this case the value of $c_{\mathcal{F}_{i}}$ is controlled to stay at least at a distance $d^{t}_{u}$ from the boundary $\partial\mathcal{F}_{i}$. We indicate with $\mathcal{Z}_{i}$ the region set where the $i$–th robot is allowed to go in the next time step. Notice that the dimension of the region $\mathcal{Z}_{i}$ can be controlled by $d_{u}^{t}$. To be conservative, the value of $d_{u}^{t}$ should be determined experimentally as the maximum tracking error, and it can be set as a function of the velocity. In Figure 5b, we show the case where we have measurement uncertainties. We illustrate both the $i$–th and the $j$–th robots’ positions and the ellipse of uncertainty $\mathcal{E}_{i,j}$. In this case, $\mathcal{K}_{i}$ indicates the region where the centroid $c_{\mathcal{F}_{i}}$ has to be constrained to preserve safety. By relying on geometric reasoning, depending on the measurement uncertainty, when $\left(1-\frac{1}{\epsilon_{p}}\right)\|p_{i}-p_{j}\|<k\sqrt{\lambda^{\max}}+\delta$, we can tune the parameter $d_{u}^{m}$ as
$d_{u}^{m}\geq\left|\left(1-\frac{1}{\epsilon_{p}}\right)\|p_{i}-p_{j}\|-\left({k\sqrt{\lambda^{\max}}}+\delta\right)\right|+\delta,$ (14)
where $\lambda^{\max}$ is the maximum eigenvalue of the covariance matrix associated with the measurement and $k$ is a scale factor that defines the confidence ellipse (e.g., $k\approx 2.795$ for a confidence level of $95\%$).
From a practical point of view, the value of $d_{u}^{m}$ may be set empirically to a constant that satisfies inequality (14) in the worst-case expected scenario.
In real-world equipments the uncertainties associated with the measurements are far from being constant values. Hence, the user may instead set $d_{u}^{m}$ dynamically, based on the current measurement uncertainties, which are reported by the equipment together with the measurement means. One method to do this, is to use the restraining method from [31]. This technique offsets the set-points of the robots from an ideal target location according to the probability distributions of relative pose observations.
On the other hand, to preserve proximity maintenance, if $\Gamma_{ij}-\|p_{i}-p_{j}\|\approx k\sqrt{\lambda^{\max}}$, we need to set $d_{u}^{m}\geq k\sqrt{\lambda^{\max}}$. Since this value may be quite high in practice, it is sufficient to impose that if the mean value of the neighbor position is out of the proximity area, then we can re-project its position to the closest point inside the proximity area. In this way, we can effectively maintain the proximity with the neighbors.
Finally, we show the combination of tracking error and measurement uncertainties, where $d_{u}=d_{u}^{m}+d_{u}^{t}$ in Figure 5c.
Figure 5: Dealing with uncertainties: (a) tracking error case, (b) measurement uncertainty case, (c) combined tracking error and measurement uncertainty. The robots are indicated with the cyan circles. We represent in light yellow the $\mathcal{F}_{i}$ set, in dark yellow the set $\mathcal{K}_{i}$, in green the set $\mathcal{Z}_{i}$, in blue the ellipse of uncertainty $\mathcal{E}_{i,j}$.
Remark 1 (Tracking error and measurement uncertainties invariance property).
Since we are adopting the same strategy to account for the tracking errors and measurements uncertainty, we get the property of invariance. In practice, it does not matter how $d_{u}^{m}$ and $d_{u}^{t}$ are chosen, what matters is the sum $d_{u}$, hence in the case of underestimation of one of the two terms, it is compensated by the other.
Remark 2 (Convergence for the adaptive solution).
By relying on the adaptive solution, while safety and proximity maintenance have been taken into account, we lost the convergence guarantees. In fact, by relying on the adaptive strategy that accounts for tracking error and measurement uncertainties, a new sufficient condition is introduced, which strictly depends on $d_{u}$ and $\epsilon_{o}$. In particular, given the minimum distance between the obstacles in the environment being $d^{\min}_{o}$, in order to ensure the passage of the robot between two obstacles, it can be shown through geometrical reasoning, that we need to have $d_{u}<d^{\min}_{o}-\sqrt{2}(\delta_{i}+\delta^{o})$. Notice that, if $d^{\min}_{o}<2\sqrt{2}(\delta_{i}+\delta^{o})$, the value of $\epsilon_{o}$ also plays a role. Hence, in this case, either the value of $d_{u}$ should be selected in a less conservative way, or we need to select $\epsilon_{o}=1$. A situation where the cell is too small to allow the centroid to be at the required distance from the boundary of the cell can also occur. However, we claim that this is not an issue for safety nor for proximity maintenance, since in these cases the robot velocity approaches zero. Such situation is transitory and may occur in very crowded conditions or when the robot has to maintain proximity with two or more robots which are close to the maximum allowed distance $\Gamma_{ij}$.
IV Simulation results
In this section we show through simulations, the scalability, repeatability and robustness of our approach, and we provide a comparison with a state-of-the-art algorithm. In particular, we report a comparison with PACNav [7], an algorithm that does not rely on communication network, and shares similar features with our proposed algorithm (refer to Table I). Despite [7] performing well in relatively simple environments and with a limited number of robots, it lacks reliability in more complex scenarios. In addition, crucial requirements such as safety and proximity maintenance depend strongly on fine parameter tuning. This is a common issue in the state-of-the-art algorithms, which is overcome by our solution, where the parameter selection affects only the performance of the mission, while, safety, proximity maintenance and convergence are met if they are selected according to the theoretical findings [2]. We perform the comparison in the MRS simulator [32] in the same environment and same model dynamics for both algorithms. We simulate three scenarios with different number of robots and with different increasingly complex obstacle configurations. For each scenario, we perform $10$ runs for statistical evidence. The metrics that we took into account are:
  • The time required for all the robots to travel $42$ meters in the goal direction;
  • The success rate in terms of mutual collision avoidance and obstacle avoidance.
In Table II, we report the results obtained from PACNav [7] and from our approach. We considered the following scenario: $N=[4,9,9]$, $\delta_{i}=[0.25,0.5,0.5]$ (m), $\delta_{o}=[0.3,0.75,0.75]$ (m), and we selected the parameters as follows for all the three scenarios $r_{s,i}=4.5$ (m), $d_{1}=d_{3}=0.5$ (m), $d_{2}=d_{4}=1$ (m), $\beta_{i}^{D}=0.5$, $\Gamma_{ij}=8$ (m), $d_{u}=0.8$ (m). The adjacency matrix which defines the proximity maintenance constraints were selected as,
$A_{d}=\begin{bmatrix}0&1&0&1\\ 1&0&1&0\\ 0&1&0&1\\ 1&0&1&0\end{bmatrix},$
and
$A_{d}=\begin{bmatrix}0&1&0&1&0&0&0&0&0\\ 1&0&1&0&1&0&0&0&0\\ 0&1&0&0&0&1&0&0&0\\ 1&0&0&0&1&0&1&0&0\\ 0&1&0&1&0&1&0&1&0\\ 0&0&1&0&1&0&0&0&1\\ 0&0&0&1&0&0&0&1&0\\ 0&0&0&0&1&0&1&0&1\\ 0&0&0&0&0&1&0&1&0\\ \end{bmatrix}.$
For our approach we considered two different values for the parameters $\epsilon=\epsilon_{p}=\epsilon_{o}$. The parameter selection for PACNav [7] has been done by trial and error, on the basis of the environment and the number of robots. Besides that, as seen in from Table II, the approach does not scale in terms of performance (convergence, safety, and proximity maintenance) when the number of robots grows, due to the complex superimposition of forces given by the surroundings.
PACNav [7] Our ($\epsilon=2.00$) Our ($\epsilon=1.052$)
Scenario avg. time (s) SR avg. time (s) SR avg. time (s) SR
1 55.01 1.00 39.96 1.00 24.23 1.00
2 - 0.00 45.71 1.00 29.53 1.00
3 - 0.00 120.33 1.00 55.87 1.00
Table II: The simulation results obtained in terms of time to reach the goal and success rate for $10$ instances in $3$ different scenarios, comparing PACNav [7] with our method in two different parameter configurations.
V Experimental Results
To validate our algorithm, we implemented it on real robotic platforms. We tested the proposed approach in two different scenarios: in the field with artificial obstacles, and in the forest. While in the former case we designed the distances between the obstacles and we relied on GNSS signal, in the forest we faced more challenging conditions both in terms of obstacle density, and in terms of localization, since we did not rely on GNSS signal. We started in a more controlled environment in order to have easier access to a ground truth measurement of the positions of the robots to evaluate the algorithm, and also for a higher repeatability of the experiment, because of the reduced complexity of the obstacle distribution.
Obstacle detection
To detect the surrounding obstacles we relied on the RPLiDAR A3. In order to have a robust and safer perception system, we introduced a double layer of obstacles. The first layer of obstacles is given by the occupancy map built online by Hector [33], while the second layer is given by the raw pointcloud provided by the RPLiDAR A3. Given this information, we cluster the points on the basis of their spatial distribution. Then we compute the closest point to the robot position for each cluster, in order to estimate the position of the obstacles in the scene. By doing that, we exploit both the reactiveness of the raw data coming from the LiDAR pointcloud, and at the same time, the scan-matching based obstacle representation, which is provided by the Hector mapping algorithm.
Relative localization
The relative positions of the neighboring robots are obtained by relying on the UltraViolet Direction And Ranging (UVDAR) technology [34]. The sensor system operates by transmission of modulated ultraviolet light from active on-board markers. These markers consist of blinking ultraviolet LEDs and their signal is retrieved by on-board cameras with tailored filters. Each of these markers is individually identifiable according to its blinking signal, and retrieving a set of them corresponding to a specific Unmanned Aerial Vehicle (UAV) allows the system to estimate the UAV’s relative pose.
V-A Experiments in a field
We tested our approach on $N=3$ aerial robots in the field with artificial obstacles. We considered $\delta_{i}=1$ (m), $\delta^{o}=0.05$ (m), $r_{s,i}=6$ (m), $d_{1}=d_{3}=1$ (m), $d_{2}=d_{4}=1$ (m), $\beta_{i}^{D}=2.0$, $\Gamma_{ij}=12$ (m), $d_{u}=1.2$ (m), $\epsilon_{p}=\epsilon_{o}=2$, and the adjacency matrix that defines the flocking constraints is $A_{d}=\left[\begin{smallmatrix}0&0&1\\ 0&0&1\\ 1&1&0\end{smallmatrix}\right]$. In this first experiment, we positioned $N_{o}=7$ artificial obstacles that mimic a portion of the forest. The robots relied on GNSS. In accordance with theoretical findings, despite the presence of uncertainties in the perception, and tracking errors due to model mismatch and external disturbances, each robot safely converged to the goal regions maintaining the desired flocking constraints.
In Figure 6, we depict five snapshots at different time instants during the mission. The robots are depicted with colored circles depending on the time step, in transparency we depict $\delta_{i}+d_{u}$, which represents the amount of uncertainties that we are considering. The obstacles are depicted in black. The green crosses indicate the goal positions. As seen, the robots respected the flocking and the safety constraints and converged successfully to the goal regions.
Figure 6: Experimental result with $N=3$ aerial robots. We depicted $5$ snapshots in different time instants. Green crosses indicate the goal positions and the black circles indicate the obstacles. Transparent areas have radius equal to $d_{u}+\delta_{i}$.
Figure 7: On-board relative localization of UAV 1 and UAV 3 from the viewpoint of the UAV 2.
Figure 8: Distances measured by UVDAR system on UAV 2 from UAV 1 (blue) and UAV 3 (red). The flocking and safety constraints are denoted in magenta.
In Figures 7 and 8, we show the positions of the neighboring robots from the perspective of UAV $3$. The data provided by the onboard sensors is quite noisy, especially when distances between robots are significant and/or when occlusions due to the obstacles or other robots occur. Nevertheless, from Figure 8, it is clear that safety constraints are respected. On the other hand, the flocking constraint appears to exceed the limit of $\Gamma_{ij}=12$ (m) at certain times. By plotting the distances from the odometry data (Figure 9), it is clear that these are actually jumps in the measurements and not in the actual relative positions.
Figure 9: Distances between UAVs using odometry.
When this situation occurs, the uncertainty of the measure can be quite high, and since we set $d_{u}=1.2$ (m) as a constant value it may cause an appearance of constraint violation. From a practical point of view, the neighbor position is set to be at a maximum distance of $\Gamma_{ij}-d_{u}$, in order to maintain a well-defined cell, or we simply discard that value if the associated uncertainty is unreasonably large. Another solution would be to increase the value of $d_{u}$, but this would be at the cost of a drop of the performance. Similar evaluation was done for UAV $3$ (see Figure 1011). In this case the flocking constraints are active on both neighbors and, as it seen in Figure 9, the constraints are satisfied.
Figure 10: On-board relative localization of UAV 1 and UAV 2 from UAV 3 viewpoint.
Figure 11: Distances of UAV 1 (red) and UAV 2 (blue) measured by UVDAR system on UAV 3. The flocking and safety constraints are denoted in magenta.
V-B Experiments in a dense environment
The second experimental deployment is conducted in the forest, where we do not have access to the GNSS signal, nor we can set a priori the obstacle positions. In this case, we change parameters to deal with more crowded conditions, according to the theoretical findings. The goal positions are set to be $40$ (m) ahead with respect the initial position of each robot, notice that we do not need to have a common reference frame, however we need to satisfy the requirement stated in Corollary 1, i.e., $\Gamma_{ij}>\|e_{i}-e_{j}\|>\Delta_{ij}$, $\forall i,j=1\dots N$ with $i\neq j$. We tested the algorithm with up to $N=4$, by selecting $\delta_{i}=0.3$ (m), $\delta^{o}=0.05$ (m), $r_{s,i}=6$ (m), $d_{1}=d_{3}=0.5$ (m), $d_{2}=d_{4}=1$ (m), $\beta_{i}^{D}=2.5$, $\Gamma_{ij}=12$ (m), $d_{u}=0.8$ (m), $\epsilon_{p}=\epsilon_{o}=1.052$, and the adjacency matrix that defines the flocking constraints is
$A_{d}=\begin{bmatrix}0&1&0&1\\ 1&0&1&0\\ 0&1&0&1\\ 1&0&1&0\end{bmatrix}.$
In Figure 12, we show the trajectory of UAV $2$, the positions of the other robots measured by UVDAR system, and the map built by UAV $2$ by using Hector SLAM during the mission. In Figure 13, we depict the distances between UAV $2$ and the other robots according to the UVDAR measurements. Similar considerations as in the previous experiments apply also in this case. In Table III, we report the average velocities, the maximum velocities, and the time to reach the goal for each UAV during the mission. To verify the success of the mission, we report in Figure 15 the actual trajectories for all the 4 robots in the mission, while in Figure 14 we show the distances between the UAVs during the mission. As we expected the safety constrained and the flocking constraints (according to the adjancency matrix $A_{d}$) are satisfied.
To show reliability and robustness to parameter changes, we report in Table IV the results obtained from other four experiments in the forest, with $N=3$ UAVs. Since the system is intrinsically chaotic it does not make sense to compare the results obtained, even if, as we can expect, a relation between $\beta^{D}$ and the time to reach the goal can be inspected from Table IV. Nevertheless, through the table we want to show the repeatability of the experiments and the robustness to parameter changes. The parameters for each experiment were set as follows: ##\##experiment $=[1,2,3,4]$, $N=[3,3,3,3]$, $\delta_{i}=[0.3,0.3,0.3,0.3]$ (m), $\delta^{o}=[0.05,0.05,0.05,0.05]$ (m) $r_{s,i}=[6,6,6,6]$ (m), $d_{1}=d_{3}=[0.5,0.5,0.5,0.5]$ (m), $d_{2}=d_{4}=[1,1,1,1]$ (m), $\beta_{i}^{D}=[2.0,2.0,1.0,0.6]$, $\Gamma_{ij}=[12,10,10,10]$ (m), $d_{u}=[0.75,0.75,0.75,0.75]$ (m), $\epsilon_{p}=\epsilon_{o}=[1.052,1.052,1.010,1.052]$, and the adjacency matrix that defines the flocking constraints is the same for all the four experiments, $A_{d}=\left[\begin{smallmatrix}0&0&1\\ 0&0&1\\ 1&1&0\end{smallmatrix}\right].$
The videos of the experiments discussed in this section can be found in the multimedia material.
UAV UAV1 UAV2 UAV3 UAV4
avg. speed (m/s) 0.361 0.329 0.373 0.370
max. speed (m/s) 1.256 1.224 1.427 1.095
time (s) 150.61 174.95 159.20 174.20
Table III: Average speed, maximum speed, and time to reach the goal of each UAV in the experiment from Figure 12.
Figure 12: Experiment in the forest. The trajectory of UAV $2$ is depicted in green. The positions measured through UVDAR by UAV $2$ are depicted in orange (UAV $1$), blue (UAV $3$), and red (UAV $4$). The map is built by UAV $1$ by means of Hector mapping is also represented, in particular the free space is white, obstacles are black and the unexplored area is grey.
Figure 13: Distances measured by UVDAR system on UAV 2 from UAV 1 (blue), UAV 3 (orange), and UAV 4 (red). The flocking and safety constraints are in magenta .
Figure 14: Distances between UAVs using odometry.
Figure 15: Experiment in the forest. The trajectory of UAV $1$ is depicted in orange, UAV $2$ in green, UAV $3$ in blue, and UAV $4$ in red. The map built by UAV $2$ by means of Hector mapping is also represented, in particular the free space is white, obstacles are black and the unexplored area is grey.
UAV1 UAV2 UAV3
Exp. 1 ($\beta^{D}=2.0$, $\epsilon_{o(p)}=1.052$, $\Gamma_{ij}=12$)
avg. speed (m/s) 0.309 0.249 0.263
max. speed (m/s) 1.210 1.255 1.226
time (s) 278.43 280.54 245.12
Exp. 2 ($\beta^{D}=2.0$, $\epsilon_{o(p)}=1.052$, $\Gamma_{ij}=10$)
avg. speed (m/s) 0.447 0.415 0.382
max. speed (m/s) 1.150 1.170 1.230
time (s) 148.00 163.81 159.91
Exp. 3 ($\beta^{D}=1.0$, $\epsilon_{o(p)}=1.010$, $\Gamma_{ij}=10$)
avg. speed (m/s) 0.661 0.497 0.740
max. speed (m/s) 1.229 1.245 1.374
time (s) 85.16 117.16 79.20
Exp. 4 ($\beta^{D}=0.6$, $\epsilon_{o(p)}=1.052$, $\Gamma_{ij}=10$)
avg. speed (m/s) 0.605 0.651 0.794
max. speed (m/s) 1.175 1.281 1.310
time (s) 82.49 80.33 62.30
Table IV: Comparison of average speed, maximum speed, and time to reach the goal in four different experiments in the forest, showing repeatability and robustness to parameters changes.
VI Conclusions
In this paper, we presented a novel distributed algorithm to swarm reliably and effectively in dense cluttered environments such as a forest, accounting for tracking errors and measurements uncertainties. We provide sufficient conditions for the theoretical guarantees of safety, proximity maintenance and convergence, compare it with state of the art algorithms, and finally we validate our solution through experiments in the field and in dense environment such as a forest. Our approach presents some imitations; the algorithm is limited to work on a plane, i.e., the reference in the altitude for each robot is constant and equal. To ensure the convergence towards the goal region, we assume that the obstacles are at a certain distance one each other (see Corollary 1). When these conditions are not met, providing to the algorithm an higher level re-planner can be beneficial to enhance performance. Future research directions will be focused on an effective 3D extension of the Lloyd algorithm and to increasing the agility of the swarm.
CNN
Convolutional Neural Network
IR
infrared
GNSS
Global Navigation Satellite System
mo-cap
Motion capture
MPC
Model Predictive Control
MRS
Multi-robot Systems Group
ML
Machine Learning
MAV
Micro-scale Unmanned Aerial Vehicle
UAV
Unmanned Aerial Vehicle
UV
ultraviolet
UVDAR
UltraViolet Direction And Ranging
UT
Unscented Transform
RTK
Real-Time Kinematic
ROS
Robot Operating System
w.r.t.
with respect to
FEC
formation-enforcing control
DIFEC
distributed formation-enforcing control
LiDAR
Light Detection And Ranging
UWB
Ultra-wideband

### v2

@@
-Swarming in the Wild: A Distributed Communication-less Lloyd-based Algorithm dealing with Uncertainties
+Distributed Lloyd-Based Algorithm for Uncertainty-Aware Multi-Robot Under-Canopy Flocking
 Abstract
-In this work, we present a distributed algorithm for swarming in complex environments that operates with no communication, no a priori information about the environment, and using only onboard sensing and computation capabilities.
-We provide sufficient conditions to guarantee that each robot reaches its goal region in a finite time, avoiding collisions with obstacles and other robots without exceeding a desired maximum distance from a predefined set of neighbors (flocking constraint).
-In addition, we show how the proposed algorithm can deal with tracking errors and onboard sensing errors without violating safety and proximity constraints, still providing the conditions for having convergence towards the goal region.
-To validate the approach, we provide experiments in the field.
-We tested our algorithm in GNSS-denied environments i.e., a dense forest, where fully autonomous aerial robots swarmed safely to the desired destinations, by relying only on onboard sensors, i.e., without a communication network.
-This work marks the initial deployment of a fully distributed system where there is no communication between the robots, nor reliance on any global localization system, which at the same time it ensures safety and convergence towards the goal within such complex environments.
+In this letter, we present a distributed algorithm for flocking in complex environments that operates at constant altitude, without explicit communication, no a priori information about the environment, and by using only on-board sensing and computation capabilities.
+We provide sufficient conditions to guarantee that each robot reaches its goal region in a finite time, avoiding collisions with obstacles and other robots without exceeding a desired maximum distance from a predefined set of neighbors (flocking or proximity constraint).
+The proposed approach allows to operate in crowded scenarios and to deal with tracking errors and on-board sensing errors, without violating safety and proximity constraints.
+The algorithm was verified through simulations with varying number of UAVs and also through numerous real-world experiments in a dense forest involving up to four UAVs.
 Index Terms:
@@
 The definition of flocking is provided by the seminal work of Reynolds [1], which defines it as a non-colliding aggregate motion.
-Depending on the environments where robots operate, specific factors like the reliability or availability of direct communication between robots may be compromised.
+Depending on the environments where robots operate, the reliability or availability of explicit communication between robots may be compromised.
 In this work, we propose a Lloyd-based algorithm that relies only on on-board sensors.
-We aim to synthesize a distributed control algorithm, where each robot safely reaches the desired position, maintaining proximity between robots in the team.
-In particular, we provide sufficient conditions to guarantee safety, proximity maintenance and convergence towards the goal in cluttered environments by accounting for bounded tracking errors and uncertainties in the estimation of neighbors and obstacles positions.
-Our algorithm takes inspiration from nature; in fact, just like flock of birds or a school of fish, the robots do not need a leader and each robot makes its own decisions based on relative information, without relying on explicit communication.
-We validate our approach through an extensive simulation and experimental campaign in the field.
-In Figure 1, we illustrate a snapshot during a mission in the forest.
-This work represents the first deployment of a fully distributed system with absence of any communication between the robots and any global localization system, which also provides conditions for safety and convergence towards the goal in such a complex environments.
-Figure 1: Deployment of the proposed algorithm in a real-world scenario with $3$ UAVs navigating in a dense forest.
+We aim to synthesize a distributed algorithm, where each robot safely reaches the desired position, while maintaining proximity between robots in the team.
+We provide sufficient conditions to guarantee safety, proximity maintenance and convergence towards the goal in cluttered environments by accounting for bounded tracking errors and uncertainties in the estimation of neighbors and obstacles’ positions.
+Our algorithm takes inspiration from the nature; just like flock of birds or a school of fish, the robots do not need a leader and each robot makes its own decisions based on relative information, without relying on explicit communication.
+We validate our approach through simulations and experiments in the forest (see Figure 1).
+Figure 1: Deployment of the proposed algorithm in a real-world scenario with 33 UAVs navigating in a forest.
 I-A Related works
 Prior literature provides a considerable amount of algorithms for distributed motion coordination of multiple robots, but only a few of them can manage to work in a cluttered environment by also imposing the flocking constraint.
-Among them, a smaller subset is devoted to applications in the field without relying on direct communication between the robots.
-To position our algorithm within the state of the art, we considered the following features:
-  • 1.
-    Distributed control: each robot makes decisions on the basis of local information, which is obtained only by relying on onboard sensors;
-  • 2.
-    Asynchronous control: each robot makes decisions independently, in asynchronous fashion;
-  • 3.
-    Communication independence: the robots do not need to communicate, since the algorithm does not rely on unobservable information, e.g., future trajectories, velocities, and accelerations;
-  • 4.
-    Safety: the system is provably safe, i.e., given the upper bounds on uncertainties (see item 7), collision and obstacle avoidance is guaranteed at any time;
-  • 5.
-    Convergence: the sufficient conditions for the robots to converge to their goal locations are provided (i.e., absence of deadlock and live-lock);
-  • 6.
-    Proximity maintenance: the algorithm allows to set hard constraints on the maximum distance between robots;
-  • 7.
-    Uncertainties: the system is provably safe and respects proximity constraints even in the presence of bounded tracking errors and/or bounded measurement uncertainties;
-  • 8.
-    Cluttered and highly dense environments: the algorithm can safely manage the presence of static obstacles in the environment.
-It can also handle highly dense cluttered environments, such as a dense forest.
-  • 9.
-    Onboard localization: the algorithm is tested by relying on onboard localization in unknown environments.
-  • 10.
-    Experiments: the proposed algorithm is tested on real robotic platforms in unstructured environments;
-  • 11.
-    Open source code: the algorithm is open source on Github.
+Among them, a smaller subset is devoted to applications in the field without relying on explicit communication between the robots.
+To position our algorithm within the state of the art, we consider the following features.
+Distributed control: each robot makes decisions on the basis of local information, which is obtained by relying only on on-board sensors.
+Asynchronous control: each robot makes decisions independently, in asynchronous fashion.
+Without explicit communication: the robots do not exchange time-dependent information such as future trajectories, current positions or velocities.
+Moreover, they do not rely on a common reference frame.
+Safety: collision and obstacle avoidance is guaranteed at any time.
+Convergence: the sufficient conditions for the robots to converge to their goal locations are provided.
+Proximity maintenance: the distance between robots is guaranteed to be below the set threshold on maximum distance.
+Uncertainties: the system is provably safe and respects proximity constraints even in the presence of bounded tracking errors and/or bounded measurement uncertainties.
+Cluttered and highly dense environments: the algorithm can operate in cluttered environments such as a forest.
+In particular, in absence of uncertainties and convergence guarantees, the robots can operate in environments where the minimum distance between two obstacles is large enough to allow the single robot to pass.
+To ensure convergence and account for uncertainties, a sufficient condition is that the obstacles are spaced far enough to allow passage of twice the robot’s size.
+On-board localization: the algorithm is verified in unknown environments relying only on on-board localization.
+Experiments: the algorithm is tested on real robotic platforms in unstructured environments.
+Open source code: the algorithm is open source.
 Method [2] [3] [4] [5] [6] [7] [8] [9] [10] [11] [12] [13] [14] [15] [16] [17] [18] [19] [20] [21] [22] [23] [24] [25] [26] ours
@@
 Open Source Code $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
-Table I: List of state of the art algorithms. We indicate with “$\times$” if the algorithm does not implement the feature, “✓” if it does, and “o” if it does only partially.
+Table I: List of state-of-the-art algorithms. We indicate with $\times$” if the algorithm does not implement the feature as described in Section I-A, “✓” if it does, and “o” if it does only partially.
 In Table I, we report an overview of the most closely related algorithms in the literature.
-We mark with “$\times$” if the algorithm does not implement the feature, “✓” if it does, and finally, “o” if it does but only partially.
-As shown in Table I, our algorithm is unmatched in fulfilling all the desired features of interest.
+We mark with $\times$” if the algorithm does not implement the feature, “✓” if it does, and finally, “o” if it does but only partially.
+As shown in Table I, our algorithm is unmatched in fulfilling all the specified features of interest.
 In fact, despite [5, 4] proposing valid solutions to the problem, their main limitations are the requirement of a trajectory broadcasting network, the absence of both convergence analysis and a rigorous account of tracking errors and uncertainties.
 Similarly, in [6] the robots need to broadcast information such as poses, velocities, poses covariance, and global extrinsic transformations.
-On the other hand, [7, 8], do not rely on a communication network, but they exhibit degraded performance in crowded scenarios as well as they lack of safety, proximity maintenance, convergence guarantees, and robustness to tracking errors and uncertainties.
-In [10], the authors present a distributed solution, which in theory does not rely on communication and accounts for uncertainties.
+On the other hand, [7, 8], do not rely on an explicit communication network, but they exhibit degraded performance in crowded scenarios, lack of safety, proximity maintenance, convergence guarantees analysis, and robustness to tracking errors and uncertainties.
+In [10], the authors present a distributed solution, which in theory does not rely on explicit communication and accounts for uncertainties.
 However, in the experiments they rely on motion capture measurements adding limited noise errors, which is reasonable only in structured environment (i.e., laboratory conditions).
-In Table  I, we report other state-of-the-art algorithms.
-Notably, all of them share limitations related to the use of a communication network and/or to not rigorously account for safety, convergence, proximity maintenance and the uncertainties.
-Conversely in our work, we provide a method which implements all the fore-mentioned features as well as a clear strategy for the selection of the relevant parameters, which in the state-of-the-art is often an unclear or empirical process.
+With respect to [3, 27], where the authors introduced the flocking behavior in the Lloyd-based framework, we do not impose the line-of-sight constraints, i.e., two robots are considered in the proximity if they are close enough.
+This allows to increase the robots mobility and also to not rely on explicit communication.
+In Table  I, we report additional relevant state-of-the-art algorithms.
+Notably, all of them share limitations related to the use of an explicit communication network and/or to not rigorously account for safety, convergence, proximity maintenance and the uncertainties.
+Conversely, in our work, we provide a method which implements all the fore-mentioned features as well as a comprehensive rule for the selection of the algorithm’s parameters based on measurable quantities (e.g., tracking error, or covariance of state estimates), which is often an unclear or empirical process in the state-of-the-art algorithms.
 I-B Contribution and paper organization
-Our contribution is fivefold:
-  • We propose a novel algorithm for flocking control in dense cluttered environments exploiting some of the theoretical findings in [2].
-With respect to [3], where the authors introduced the flocking behavior in the Lloyd-based framework, we do not impose the line-of-sight constraint, i.e., two robots are considered in the proximity if they are close enough.
-This allows to increase the robots mobility and also to not rely on communication.
-At the same time, it introduces high uncertainties in the position estimation of the neighbors when robots are not in the line of sight.
-  • We provide sufficient conditions for safety, proximity maintenance, and convergence even in cluttered environments and with flocking constraints.
-  • We propose a novel adaptive method to account for tracking errors and measurement uncertainties.
-  • We tailor the use of Convex Weighted Voronoi Diagrams (CWVD) [13] for our problem.
-By doing that, we considerably improve the performance of the algorithm in dense cluttered environments such as a dense forest.
-  • We validate the algorithm in the forest, relying only on on-board sensors i.e., IMU, 2-D Lidar and UVDAR system, without relying neither on a communication network nor GNSS signal.
-The paper is organized as follows: Sec.
-II provides the problem description, in Sec.
-III we present the background material and the proposed approach, in Sec.
-IV we provide simulation results and a comparison with a state-of-the-art approach.
-In Sec.
-V we validate the proposed approach through experimental results in the field, and finally in Sec.
-VI we draw the conclusions and the future research directions.
+Our contribution is threefold.
+First, we propose a novel algorithm for flocking control in dense cluttered environments.
+The key novelties lie in our use of an adaptive control strategy to compensate for tracking errors and measurement uncertainties, combined with Convex Weighted Voronoi Diagrams (CWVD) to enhance the performance in a dense forest scenarios.
+Second, we provide sufficient conditions for safety, proximity maintenance, and convergence in cluttered environments and in the presence of flocking constraints.
+Finally, we validate the algorithm in the forest, relying only on on-board sensors, i.e., Inertial Measurement Unit (IMU), 2-D Lidar and UVDAR system [28], without relying neither on an explicit communication network nor GNSS signal.
 II Problem description
 Let us denote the positions of the robots with $p=[p_{1},\dots,p_{N}]^{\top}$, where $p_{i}=[x_{i},y_{i}]^{\top}$ and the obstacle positions by $o=[o_{1},\dots,o_{N_{o}}]^{\top}$, $o_{k}=[x_{k},y_{k}]^{\top}$.
-The mission space is represented by $\mathcal{Q}$, the weighting function that assess the weights of the points $q\in\mathcal{Q}$ is $\varphi_{i}(q):\mathcal{Q}\rightarrow\mathbb{R}_{+}$, while $\delta_{i}$ and $\delta_{k}^{o}$ are the radius of encumbrance of the $i$–th robot and the $k$–th obstacle respectively.
+The mission space is represented by $\mathcal{Q}\in\mathbb{R}^{2}$, the weighting function that assesses the weights of the points $q\in\mathcal{Q}$ is $\varphi_{i}(q):\mathcal{Q}\rightarrow\mathbb{R}_{+}$, while $\delta_{i}$ and $\delta_{k}^{o}$ are the occupancy radius of the ii–th robot and the kk–th obstacle respectively.
 The problem that we aim to solve reads as follow:
@@
 Let $N$ be the number of robots in the swarm, let $\mathcal{N}_{i}$ be the set of neighbors of the robot $i$, and $\bar{\mathcal{N}}_{i}\subseteq\mathcal{N}_{i}$ the set of neighbors that has to remain connected the entire mission.
-We want to synthesize a distributed control law, such that the swarm maintains the desired proximity constraints, which is characterized by a set of maximum distances $\Gamma_{ij},\,\,\,\forall i,\forall j\in\bar{\mathcal{N}}_{i}$.
+We want to synthesize a distributed control law, such that the swarm maintains the desired proximity constraints, which is characterized by a set of maximum distances $\Gamma_{ij}=\Gamma_{ji},\,\,\,\forall i\in{1,\dots,N},\forall j\in\bar{\mathcal{N}}_{i}$.
 Each robot in the swarm has to reach its goal region avoiding collisions with other robots and obstacles.
-Assumption 1.
-Each robot $i$ can perceive the relative position of robot (obstacle) $j$ if $j\in\mathcal{N}_{i}$ ($j\in\mathcal{O}_{i}$).
-Where $j\in\mathcal{N}_{i}$ if $\|p_{i}-p_{j}\|\leq R_{s}$ ($j\in\mathcal{O}_{i}$ if $\|p_{i}-o_{j}\|\leq R_{s}$), with $R_{s}$ indicating the sensing range, equal for all the robots.
-Assumption 2.
-The flocking constraints are symmetric, i.e., $\|p_{i}-p_{j}\|\leq\Gamma_{ij},\forall j\in\bar{\mathcal{N}}_{i},\forall i=1,\dots,N$, where $\Gamma_{ij}=\Gamma_{ji}<R_{s}$.
-Assumption 3.
-The size of the neighboring robots and obstacles is known and it can be approximated with circles/cylinders.
-Assumption 4.
-In our theoretical analysis, we assume that the robots are holonomic agents.
-Assumption 5.
-From Section III-C, we assume more complex dynamical systems and a more realistic perception system, where the tracking error is bounded as well as the error on the measured distances from neighbors and obstacles.
 III Proposed Approach
-The presented approach is grounded in the Lloyd algorithm [27]. In the literature, its distributed nature was exploited to develop novel algorithms for several collective behaviors [3, 2, 13, 28, 29].
-Let us consider the following cost function
-[…]
-where $\mathcal{V}_{i}=\{\mathcal{V}_{i}^{p}\cap\mathcal{V}_{i}^{o}\}$ with $\mathcal{V}_{i}^{p}$ being the $i$–th Convex Weighted Voronoi Diagram (CWVD) introduced in [13], which accounts only for neighboring robots positions
+The presented approach is grounded in the Lloyd algorithm [29]. Let us consider the following cost function
+$J_{\operatorname{cov}}({p},\mathcal{V})=\sum_{i\in\mathcal{N}_{i}}\int_{\mathcal{V}_{i}}\left\|q-p_{i}\right\|^{2}\varphi_{i}(q)dq,$ (1)
+where $\mathcal{V}=\{\mathcal{V}_{1},\dots,\mathcal{V}_{N}\}$, $\mathcal{V}_{i}=\mathcal{V}_{i}^{p}\cap\mathcal{V}_{i}^{o}$ with $\mathcal{V}_{i}^{p}$ being the ii–th Convex Weighted Voronoi Diagram (CWVD) introduced in [13], which accounts only for neighboring robots’ positions
 $\small\begin{split}\mathcal{V}_{i}^{p}=\{q\in\mathcal{Q}\mid w_{i}^{x}\cos\alpha_{ij}+w_{i}^{y}\sin\alpha_{ij}<\frac{1}{\epsilon_{p}}&\left\|p_{i}-p_{j}\right\|,\\ &\forall j\in\mathcal{N}_{i}\}.\end{split}$ (2)
-While $\mathcal{V}_{i}^{o}$ is the $i$–th CWVD accounting only for neighboring obstacles positions
+While $\mathcal{V}_{i}^{o}$ is the ii–th CWVD accounting only for neighboring obstacles positions
 $\small\begin{split}\mathcal{V}_{i}^{o}=\{q\in\mathcal{Q}\mid w_{i}^{x}\cos\alpha_{ij}+w_{i}^{y}\sin\alpha_{ij}<\frac{1}{\epsilon_{o}}&\left\|p_{i}-o_{j}\right\|,\\ &\forall j\in\mathcal{O}_{i}\},\end{split}$ (3)
-where $w=[w^{x},w^{y}]^{\top}=q-p_{i}$ and $\alpha_{ij}=\arctan\left(\frac{y_{j}-y_{i}}{x_{j}-x_{i}}\right)$, while $\epsilon_{p(o)}\in[1,2]$, is an arbitrary real number, which regulates the cautiousness in the robot motion (i.e., $\epsilon_{p(o)}=1$ more aggressive behavior, $\epsilon_{p(o)}=2$ more cautelative behavior, notice that the latter coincides with the classic Voronoi diagram).
+where $w_{i}=[w_{i}^{x},w_{i}^{y}]^{\top}=q-p_{i}$ and $\alpha_{ij}=\arctan\left(\frac{y_{j}-y_{i}}{x_{j}-x_{i}}\right)$, while $\epsilon_{p(o)}\in[1,2]$, is a parameter, which regulates the cautiousness in the robot motion (i.e., $\epsilon_{p(o)}=1$ more aggressive behavior, $\epsilon_{p(o)}=2$ more conservative behavior, notice that the latter coincides with the classic Voronoi diagram).
 Figure 2: Representation of the Convex Weighted Voronoi Diagram (CWVD) $\mathcal{V}_{i}^{p}$ in (2) with $\epsilon_{p}=2$ and $\epsilon_{p}=1$.
@@
 In Figure 2, we depict an example of the CWVD (2) for the case where $\epsilon_{p}=2$ and $\epsilon_{p}=1$.
-By exploiting Assumption 4, we can impose $\dot{p}_{i}=u_{i}$. By following the gradient descent of the cost function (1), we obtain the following proportional control law
+By following the gradient descent of the cost function (1), we obtain the following proportional control law
 $\dot{p}_{i}\left(\mathcal{V}_{i}\right)=-k_{p,i}\left(p_{i}-c_{\mathcal{V}_{i}}\right),$ (4)
 where $k_{p,i}>0$ is a positive real number, and
+$c_{\mathcal{V}_{i}}=\frac{\int_{\mathcal{V}_{i}}q\varphi_{i}(q)dq}{\int_{\mathcal{V}_{i}}\varphi_{i}(q)dq}$ (5)
+is the centroid position evaluated over the $\mathcal{V}_{i}$ cell.
+By imposing (4), assuming a time-invariant $\varphi_{i}(q)$, each robot $i\in 1,\dots,N$, converges asymptotically to the centroid position $c_{\mathcal{V}_{i}}$ [30].
+In the following, we show how we can attain a safe and effective algorithm for multi-robot swarming by introducing suitable modifications to the geometry of cell $\mathcal{V}_{i}$ and crafting an appropriate weighting function $\varphi_{i}(q)$
+III-A Reshaping the cell geometry $\mathcal{V}_{i}$
+The reshape of the cell geometry is both necessary and sufficient to ensure safety guarantees.
+Hence, we define the cell $\mathcal{A}_{i}=\tilde{\mathcal{V}}_{i}^{p}\cap\tilde{\mathcal{V}}_{i}^{o}\cap\mathcal{S}_{i}$, where
 […]
-is the centroid position evaluated over the $\mathcal{V}_{i}$ cell.
-By imposing (4), assuming a time-invariant $\varphi_{i}(q)$, each robot $i$, $\forall i,1\dots,N$, converges asymptotically to the centroid position $c_{\mathcal{V}_{i}}$ [30].
-In the following, we show how, by introducing suitable modifications to the geometry of cell $\mathcal{V}_{i}$ and crafting an appropriate weighting function $\varphi_{i}(q)$, we can attain a safe and effective algorithm for multi-robot swarming.
-III-A Cell geometry $\mathcal{F}_{i}$
-The reshape of the cell geometry is both necessary and sufficient to ensure safety guarantees.
-Hence, we define the cell $\mathcal{A}_{i}=\{\tilde{\mathcal{V}}_{i}^{p}\cap\tilde{\mathcal{V}}_{i}^{o}\cap\mathcal{S}_{i}\}$, where
-$\tilde{\mathcal{V}}_{i}^{p}=\left\{\begin{array}[]{l}\left\{q\in\mathcal{V}_{i}^{p}\right\}\text{ if }\frac{1}{2}\left\|p_{i}-p_{j}\right\|\geq\Delta_{ij}\\ \left\{q\in\mathcal{Q}\mid\left\|q-p_{i}\right\|\leq\left\|q-\tilde{p}_{j}\right\|\right\}\text{ otherwise, }\end{array}\right.$ (6)
-with $j\in\mathcal{N}_{i}$, $\Delta_{ij}=\delta_{j}+\delta_{i}$, and $\tilde{p}_{j}=p_{j}+2\left(\Delta_{ij}-\frac{\left\|p_{i}-p_{j}\right\|}{2}\right)\frac{p_{i}-p_{j}}{\left\|p_{i}-p_{j}\right\|}$, takes into consideration the physical dimension of the robots.
-On the other hand,
-$\tilde{\mathcal{V}}_{i}^{o}=\left\{\begin{array}[]{l}\left\{q\in\mathcal{V}_{i}^{o}\right\}\text{ if }\frac{1}{2}\left\|p_{i}-p_{j}\right\|\geq\Delta^{o}_{ij}\\ \left\{q\in\mathcal{Q}\mid\left\|q-p_{i}\right\|\leq\left\|q-\tilde{o}_{j}\right\|\right\}\text{ otherwise, }\end{array}\right.$ (7)
-with $j\in\mathcal{O}_{i}$, $\Delta^{o}_{ij}=\delta_{i}+\delta^{o}_{j}$ and $\tilde{o}_{j}=o_{j}+2\left(\Delta^{o}_{ij}-\frac{\left\|p_{i}-o_{j}\right\|}{2}\right)\frac{p_{i}-o_{j}}{\left\|p_{i}-o_{j}\right\|}$, takes into consideration the size of the robot and the obstacles, while
+where $\Delta_{ij}=\delta_{j}+\delta_{i}$, and $\tilde{p}_{j}=p_{j}+2\left(\Delta_{ij}-\frac{\left\|p_{i}-p_{j}\right\|}{2}\right)\frac{p_{i}-p_{j}}{\left\|p_{i}-p_{j}\right\|}$, takes into consideration the physical dimension of the robots.
+To account for the obstacles,
+[…]
+where $\Delta^{o}_{ij}=\delta_{i}+\delta^{o}_{j}$ and $\tilde{o}_{j}=o_{j}+2\left(\Delta^{o}_{ij}-\frac{\left\|p_{i}-o_{j}\right\|}{2}\right)\frac{p_{i}-o_{j}}{\left\|p_{i}-o_{j}\right\|}$, takes into consideration the size of the robot and the obstacles, while
 $\mathcal{S}_{i}=\left\{q\in\mathcal{Q}\mid\left\|q-p_{i}\right\|\leq r_{s,i}\right\},$ (8)
-with $r_{s,i}$ is half of the sensing radius $R_{s}$.
-Theorem 1 (Safety [2]).
-By imposing the control $\dot{p}_{i}(\mathcal{A}_{i})$ in (4), collision avoidance is guaranteed at every instant of time.
+where $r_{s,i}$ is half of the sensing radius $R_{s}$ of the robot $i$.
+Theorem 1 (Safety).
+By imposing the control $\dot{p}_{i}(\mathcal{A}_{i})$ in (4), with $\tilde{\mathcal{V}}_{i}^{p}\cap\tilde{\mathcal{V}}_{i}^{o}\cap\mathcal{S}_{i}$, collision avoidance is guaranteed at every instant of time.
 Proof.
-The proof can be found in [2], the addition of static obstacles and CVWD does not change the result.
-In particular, by adding static obstacles and/or introducing CVWD does not alter the convexity of the set.
+The proof comes from the definition of convex set.
+Due to the fact that the set $\mathcal{A}_{i}$ is always convex, and by definition $p_{i},c_{\mathcal{A}_{i}}\in\mathcal{A}_{i}$, it follows that since the straight path from the robot position to the centroid position is a safe path, then applying $\dot{p}_{i}(\mathcal{A}_{i})$ in (4) preserves safety.
 ∎
-We further apply an additional reshaping of the cell geometry in order to maintain the proximity, i.e., to obtain swarming.
-By imposing our final cell geometry $\mathcal{F}_{i}=\{\mathcal{A}_{i}\cap\mathcal{M}_{i}\}$, where
-$\mathcal{M}_{i}=\{q\in\mathcal{Q}\mid\|q-p_{j}\|<\Gamma_{ij},\forall j\in\bar{\mathcal{N}}_{i}\},$ (9)
-we guarantee proximity maintenance while preserving safety. In Figure 3, we depict an example of cell geometry $\mathcal{F}_{i}$.
+We further apply an additional reshaping of the cell geometry in order to maintain the desired proximity of the agents, i.e., to achieve swarming behavior.
+By imposing our final cell geometry $\mathcal{F}_{i}=\mathcal{A}_{i}\cap\mathcal{M}_{i}$, where
+$\mathcal{M}_{i}=\{q\in\mathcal{Q}\mid\|q-p_{j}\|\leq\Gamma_{ij},\forall j\in\bar{\mathcal{N}}_{i}\},$ (9)
+we obtain proximity maintenance while preserving safety. In Figure 3, we depict an example of cell geometry $\mathcal{F}_{i}$.
 Figure 3: Example of $\mathcal{F}_{i}$ and $\mathcal{A}_{i}$ cell geometry with $\epsilon_{p}=\epsilon_{o}=2$.
@@
 Given that $\mathcal{F}_{i}\subseteq\mathcal{A}_{i}$, safety comes from Theorem 1.
-Since $\mathcal{F}_{i}$ is convex by definition, then the centroid $c_{\mathcal{F}_{i}}\in\mathcal{F}_{i}$.
+Since $\mathcal{M}_{i}$ is convex by definition, $\mathcal{F}_{i}$ is also convex, and thus the centroid $c_{\mathcal{F}_{i}}\in\mathcal{F}_{i}$.
 It implies that by imposing $\dot{p}_{i}(\mathcal{F}_{i})$ in (4), $p_{i}\in\mathcal{F}_{i}$ at any time.
@@
 and
-$\begin{split}\dot{\bar{p}}_{i}&=\begin{cases}-(\bar{p}_{i}-R(\pi/2-\varepsilon)e_{i})\,\,\,&\text{if}\,\,\|c_{\mathcal{A}_{i}}-p_{i}\|<d_{3}\,\land\\ &\|c_{\mathcal{A}_{i}}-c_{\mathcal{S}_{i}}\|>d_{4}\\ -(\bar{p}_{i}-e_{i})&\text{otherwise,}\end{cases}\\ \bar{p}_{i}&=e_{i}\,\,\,\text{if}\,\,\,\|p_{i}-\bar{c}_{\mathcal{A}_{i}}\|>\|p_{i}-c_{\mathcal{A}_{i}}\|\,\land\,\bar{p}_{i}=R(\pi/2-\varepsilon)e_{i}.\\ \end{split}$ (12)
-The parameters $\beta_{i}^{D},d_{1},d_{2},d_{3},d_{4}$ are positive scalars, $R(\theta)$ is the rotation matrix, $e_{i}$ is the goal position for the $i$–th robot, $\varepsilon$ is a small number, which allows to avoid rotations equal to $\pi/2$, while $\bar{{c}}_{\mathcal{A}_{i}}$ is the centroid position computed over the cell $\mathcal{A}_{i}$ with $\bar{p}_{i}\equiv e_{i}$.
-Corollary 1 (Convergence in cluttered environment and with flocking constraints).
-Let us assume we have an unbounded space cluttered with circular obstacles of radius $\delta^{o}<\Gamma_{ij}/2-2\delta$ and at a minimum distance of $\|o_{i}-o_{j}\|>4\delta+2\delta^{o}$ one to the other.
-The initial and goal positions are adequately apart from each other, i.e., $\Gamma_{ij}>\|p_{i}(0)-p_{j}(0)\|>\Delta_{ij}$, $\Gamma_{ij}>\|e_{i}-e_{j}\|>\Delta_{ij}$, $\forall i,j=1\dots N$ with $i\neq j$.
+[…]
+The parameters $\beta_{i}^{D},d_{1},d_{2},d_{3},d_{4}$ are positive scalars, $R(\theta)$ is the rotation matrix, $e_{i}$ is the goal position for the ii–th robot, $\varepsilon$ is a small number, which allows to avoid rotations equal to $\pi/2$, while $\bar{{c}}_{\mathcal{A}_{i}}$ is the centroid position computed over the cell $\mathcal{A}_{i}$ with $\bar{p}_{i}\equiv e_{i}$.
+Proposition 1 (Convergence in cluttered environment and with flocking constraints).
+Let us assume we have an unbounded space cluttered with circular obstacles of radius $\delta^{o}<\Gamma_{ij}/2-2\delta$, where $\delta=\delta_{i},\forall i=1,\dots,N$, and at a minimum distance of $\|o_{i}-o_{j}\|>4\delta+2\delta^{o}$ one to the other.
+The initial and goal positions are adequately apart from each other, i.e., $\Gamma_{ij}>\|p_{i}(0)-p_{j}(0)\|>\Delta_{ij}$, $\Gamma_{ij}\geq\|e_{i}-e_{j}\|>\Delta_{ij}$, $\forall i,j=1\dots N$ with $i\neq j$.
 By imposing the control law $\dot{p}_{i}(\mathcal{F}_{i})$ in (4), (10), (11), and (12) to compute the centroid position in (5), by properly selecting $d_{1},d_{2},d_{3},d_{4}$, and $\beta^{D}_{i}$, we have that $\lim_{t\rightarrow\infty}\|p_{i}(t)-e_{i}\|<r_{s,i},\forall i=1,\dots,N.$
-Figure 4: We depict in cyan the robots and in grey the static obstacles.
-The crosses represent the final goal position while the dashed lines the proximity constraint topology.
-In the figure we show two possible deadlock conditions: case 1.
+Figure 4: The figure shows two possible deadlock conditions: case 1.
 is due to the small inter-space between obstacles.
@@
 is due to the too short $\Gamma_{ij}$ value (or too big $\delta^{o}$ value).
+We depict in cyan the robots and in grey the static obstacles.
+The crosses represent the final goal positions while the dashed lines represent the proximity constraint topology.
 Proof.
-With respect to the proof of convergence with no obstacles and no proximity maintenance constraints, which are provided in [2], we recognize, beside the trivial initial and final conditions requirements, the addition of two potential deadlock scenarios, which are depicted in Figure 4.
+The proof provided in [2] does not consider presence of static obstacles nor proximity constraints.
+By including both of them, we recognize, beside the trivial initial and final conditions requirements, the addition of two potential deadlock scenarios, which are depicted in Figure 4.
 Notice that the addition of static obstacles is equivalent to non-cooperative static robots.
 In order to not affect the proof of convergence obtained in [2], two requirements need to be considered.
-1.
-We require a minimum distance between the obstacles of $\|o_{i}-o_{j}\|>4\delta+2\delta^{o},\,\forall i\neq j$, in order to avoid the presence of narrow hallways that can generate deadlock issues (left hand side of Figure 4).
+First, we require a minimum distance between the obstacles of $\|o_{i}-o_{j}\|>4\delta+2\delta^{o},\,\forall i\neq j$, in order to avoid the presence of narrow hallways that can generate deadlock issues (left hand side of Figure 4).
 This is a consequence of the results in [2].
@@
 Since $d_{2}>2\delta$ and $\|e_{i}-e_{j}\|>\Delta_{ij}$, then the condition $\|o_{i}-o_{j}\|>4\delta+2\delta^{o},\,\forall i\neq j$ is equivalent.
-2.
-We require that the maximum radius of the obstacles is $\delta^{o}<\Gamma_{ij}/2-2\delta$, in order to avoid strict proximity maintenance constraints (right hand side of Figure 4).
+Second, we require that the maximum radius of the obstacles is $\delta^{o}<\Gamma_{ij}/2-2\delta$, in order to avoid strict proximity maintenance constraints (right hand side of Figure 4).
 By imposing these additional conditions, the guarantees of convergence in [2] holds.
@@
 III-C Dealing with dynamics and uncertainties
-The real robot deployment requires to consider a set of additional issues, such as 1.
-More complex dynamics of the robots must be applied, i.e., $\dot{p}_{i}=u_{i}$ (Assumption 4) cannot be directly applied.
-2.
-Tracking error, which can be due to external disturbances or model mismatches, generating a difference between the desired and the actual velocity, 3.
-Measurement uncertainties, due to the limitations of the perception system providing an estimate of the positions of other robots and obstacles with a certain error.
-To account for the dynamic constraints of the robots, we synthesize the control inputs through an MPC formulation (similarly to [10, 2]), which steers each $i$–th robot towards the centroid position $c_{\mathcal{F}_{i}}$, by constraining the trajectory to not exceed its $\mathcal{F}_{i}$ cell.
+The real robot deployment requires to consider a set of additional issues.
+First, more complex dynamics of the robots must be applied as $\dot{p}_{i}=u_{i}$ is not directly applicable.
+Second, tracking error, which can be caused by external disturbances or model mismatches, generating a difference between the desired and the actual velocity, has to be taken in to account.
+Finally, we have to consider measurement uncertainties arising from limitations of the perception system providing an estimate of the positions of other robots and obstacles with a certain error.
+To account for the dynamic constraints of the robots, we synthesize the control inputs through an MPC formulation (similarly to [10, 2]), which steers each ii–th robot towards the centroid position $c_{\mathcal{F}_{i}}$, by constraining the trajectory to not exceed its $\mathcal{F}_{i}$ cell.
 By doing that, we can account for the robot dynamics without violating the safety and proximity maintenance constraints.
@@
 s.t.: ${x}_{k}=f({x}_{k-1},{u}_{k-1}),\,\,\,\forall k\in(1,N_{t}),$ (13b)
-$\,{x}_{k}\in\mathcal{X},\,{u}_{k}\in\mathcal{U},\,\,\,\forall k\in(0,N_{t})$ (13c)
-$p_{k}\in\mathcal{F},\,\,\,\forall k\in(0,N_{t})$ (13d)
+$\,{x}_{k}\in\mathcal{X},\,{u}_{k}\in\mathcal{U},\,\,\,\forall k\in(0,N_{t}),$ (13c)
+$p_{k}\in\mathcal{F},\,\,\,\forall k\in(0,N_{t}),$ (13d)
 ${x}_{0}={x}_{\textrm{init}},$ (13e)
@@
 Both the tracking errors and measurement uncertainties can be accounted for by means of an adaptive strategy.
-In fact, it can be achieved by adapting the value of $\beta_{i}$ in Equation (11) in an online fashion.
+In fact, it can be achieved by adapting the value of $\beta_{i}$ in (11) in an online fashion.
 An alternative approach would be reshaping the $\mathcal{F}_{i}$ cell, similarly to [10].
 However, this method cannot be applied in our case, since, by following this approach, the robot may fall out of its cell due to tracking error, and once it happens, both the cell and the centroid position are no longer defined.
-Conversely, the value of $\beta_{i}$ defines the aggressiveness of the $i$–th robot.
+Conversely, the value of $\beta_{i}$ defines the aggressiveness of the ii–th robot.
 A smaller value corresponds to a centroid position closer to the goal (and hence closer to the boundaries of the $\mathcal{F}_{i}$ cell, namely $\partial\mathcal{F}_{i}$), while larger values of $\beta_{i}$ push the centroid position as far as possible from the boundaries of the cell $\partial\mathcal{F}_{i}$.
-Thus, managing errors and uncertainties can be achieved by selecting a value of $\beta_{i}$ such that the centroid position stays at a controlled distance from the boundary of the cell.
+Thus, managing errors and uncertainties is achieved by selecting a value of $\beta_{i}$ such that the centroid position stays at a controlled distance from the boundary of the cell.
 This distance must be chosen in accordance with the maximum positional tracking error and the maximum error in the neighbors’ position estimates.
@@
 Figure 5a considers the tracking error; in this case the value of $c_{\mathcal{F}_{i}}$ is controlled to stay at least at a distance $d^{t}_{u}$ from the boundary $\partial\mathcal{F}_{i}$.
-We indicate with $\mathcal{Z}_{i}$ the region set where the $i$–th robot is allowed to go in the next time step.
+We indicate with $\mathcal{Z}_{i}$ the region set where the ii–th robot is allowed to go in the next time step.
 Notice that the dimension of the region $\mathcal{Z}_{i}$ can be controlled by $d_{u}^{t}$.
@@
 In Figure 5b, we show the case where we have measurement uncertainties.
-We illustrate both the $i$–th and the $j$–th robots’ positions and the ellipse of uncertainty $\mathcal{E}_{i,j}$.
+We illustrate both the ii–th and the jj–th robots’ positions and the ellipse of uncertainty $\mathcal{E}_{i,j}$.
 In this case, $\mathcal{K}_{i}$ indicates the region where the centroid $c_{\mathcal{F}_{i}}$ has to be constrained to preserve safety.
 By relying on geometric reasoning, depending on the measurement uncertainty, when $\left(1-\frac{1}{\epsilon_{p}}\right)\|p_{i}-p_{j}\|<k\sqrt{\lambda^{\max}}+\delta$, we can tune the parameter $d_{u}^{m}$ as
-$d_{u}^{m}\geq\left|\left(1-\frac{1}{\epsilon_{p}}\right)\|p_{i}-p_{j}\|-\left({k\sqrt{\lambda^{\max}}}+\delta\right)\right|+\delta,$ (14)
+$d_{u}^{m}\geq\left\|\left(1-\frac{1}{\epsilon_{p}}\right)\|p_{i}-p_{j}\|-\left({k\sqrt{\lambda^{\max}}}+\delta\right)\right\|+\delta,$ (14)
 where $\lambda^{\max}$ is the maximum eigenvalue of the covariance matrix associated with the measurement and $k$ is a scale factor that defines the confidence ellipse (e.g., $k\approx 2.795$ for a confidence level of $95\%$).
@@
 In real-world equipments the uncertainties associated with the measurements are far from being constant values.
-Hence, the user may instead set $d_{u}^{m}$ dynamically, based on the current measurement uncertainties, which are reported by the equipment together with the measurement means.
-One method to do this, is to use the restraining method from [31].
+Hence, the user can set $d_{u}^{m}$ dynamically, based on the current measurement uncertainties, which are reported by the equipment together with the measurement means.
+One method to implement this, is to use the restraining method from [31].
 This technique offsets the set-points of the robots from an ideal target location according to the probability distributions of relative pose observations.
@@
 Since this value may be quite high in practice, it is sufficient to impose that if the mean value of the neighbor position is out of the proximity area, then we can re-project its position to the closest point inside the proximity area.
-In this way, we can effectively maintain the proximity with the neighbors.
+In this way, the proximity to the neighbors is effectively maintained.
 Finally, we show the combination of tracking error and measurement uncertainties, where $d_{u}=d_{u}^{m}+d_{u}^{t}$ in Figure 5c.
@@
 By relying on the adaptive solution, while safety and proximity maintenance have been taken into account, we lost the convergence guarantees.
-In fact, by relying on the adaptive strategy that accounts for tracking error and measurement uncertainties, a new sufficient condition is introduced, which strictly depends on $d_{u}$ and $\epsilon_{o}$.
-In particular, given the minimum distance between the obstacles in the environment being $d^{\min}_{o}$, in order to ensure the passage of the robot between two obstacles, it can be shown through geometrical reasoning, that we need to have $d_{u}<d^{\min}_{o}-\sqrt{2}(\delta_{i}+\delta^{o})$.
-Notice that, if $d^{\min}_{o}<2\sqrt{2}(\delta_{i}+\delta^{o})$, the value of $\epsilon_{o}$ also plays a role.
-Hence, in this case, either the value of $d_{u}$ should be selected in a less conservative way, or we need to select $\epsilon_{o}=1$.
+In fact, by relying on the adaptive strategy that accounts for tracking error and measurement uncertainties, a new sufficient condition is introduced, which strictly depends on $d_{u}$.
+In particular, given the minimum distance between the obstacles in the environment being $d^{\min}_{o}$, in order to ensure the passage of the robot between two obstacles, it can be shown through geometrical reasoning, that we need to have $0\leq d_{u}<\frac{d^{\min}_{o}}{2}-(\delta_{i}+\delta^{o})$.
 A situation where the cell is too small to allow the centroid to be at the required distance from the boundary of the cell can also occur.
 However, we claim that this is not an issue for safety nor for proximity maintenance, since in these cases the robot velocity approaches zero.
-Such situation is transitory and may occur in very crowded conditions or when the robot has to maintain proximity with two or more robots which are close to the maximum allowed distance $\Gamma_{ij}$.
+Such situation is transitory and may occur in very crowded conditions or when the robot has to maintain proximity with two or more robots that are close to the maximum allowed distance $\Gamma_{ij}$.
 IV Simulation results
 In this section we show through simulations, the scalability, repeatability and robustness of our approach, and we provide a comparison with a state-of-the-art algorithm.
-In particular, we report a comparison with PACNav [7], an algorithm that does not rely on communication network, and shares similar features with our proposed algorithm (refer to Table I).
-Despite [7] performing well in relatively simple environments and with a limited number of robots, it lacks reliability in more complex scenarios.
+In particular, we report a comparison with PACNav [7], an algorithm that does not rely on communication network, and shares similar features with the proposed algorithm (refer to Table I).
+Notice that comparing with algorithm such as [5, 4, 6] is out of scope of this paper.
+Our method does not perform better than algorithms that can share information, such as future trajectories, among the robots.
+Despite [7] performs well in relatively simple environments and with a limited number of robots, it lacks reliability in more complex scenarios.
 In addition, crucial requirements such as safety and proximity maintenance depend strongly on fine parameter tuning.
-This is a common issue in the state-of-the-art algorithms, which is overcome by our solution, where the parameter selection affects only the performance of the mission, while, safety, proximity maintenance and convergence are met if they are selected according to the theoretical findings [2].
+This is a common issue in the state-of-the-art algorithms, which is overcome in our solution, where the parameter selection affects only the performance of the mission, while, safety, proximity maintenance and convergence are met if they are selected according to the theoretical findings.
 We perform the comparison in the MRS simulator [32] in the same environment and same model dynamics for both algorithms.
-We simulate three scenarios with different number of robots and with different increasingly complex obstacle configurations.
+We simulate three scenarios with different number of robots and with an increasingly complex obstacle configurations.
 For each scenario, we perform $10$ runs for statistical evidence.
-The metrics that we took into account are:
-  • The time required for all the robots to travel $42$ meters in the goal direction;
-  • The success rate in terms of mutual collision avoidance and obstacle avoidance.
+The metrics that we took into account are the time required for all the robots to travel $42$ meters in the goal direction and the success rate in terms of mutual collision avoidance and obstacle avoidance.
 In Table II, we report the results obtained from PACNav [7] and from our approach.
@@
 The adjacency matrix which defines the proximity maintenance constraints were selected as,
-$A_{d}=\begin{bmatrix}0&1&0&1\\ 1&0&1&0\\ 0&1&0&1\\ 1&0&1&0\end{bmatrix},$
-and
-$A_{d}=\begin{bmatrix}0&1&0&1&0&0&0&0&0\\ 1&0&1&0&1&0&0&0&0\\ 0&1&0&0&0&1&0&0&0\\ 1&0&0&0&1&0&1&0&0\\ 0&1&0&1&0&1&0&1&0\\ 0&0&1&0&1&0&0&0&1\\ 0&0&0&1&0&0&0&1&0\\ 0&0&0&0&1&0&1&0&1\\ 0&0&0&0&0&1&0&1&0\\ \end{bmatrix}.$
+$A_{d}=\left[\begin{smallmatrix}0&1&0&1\\ 1&0&1&0\\ 0&1&0&1\\ 1&0&1&0\end{smallmatrix}\right],\text{ and}\quad A_{d}=\left[\begin{smallmatrix}0&1&0&1&0&0&0&0&0\\ 1&0&1&0&1&0&0&0&0\\ 0&1&0&0&0&1&0&0&0\\ 1&0&0&0&1&0&1&0&0\\ 0&1&0&1&0&1&0&1&0\\ 0&0&1&0&1&0&0&0&1\\ 0&0&0&1&0&0&0&1&0\\ 0&0&0&0&1&0&1&0&1\\ 0&0&0&0&0&1&0&1&0\\ \end{smallmatrix}\right].$
 For our approach we considered two different values for the parameters $\epsilon=\epsilon_{p}=\epsilon_{o}$.
 The parameter selection for PACNav [7] has been done by trial and error, on the basis of the environment and the number of robots.
-Besides that, as seen in from Table II, the approach does not scale in terms of performance (convergence, safety, and proximity maintenance) when the number of robots grows, due to the complex superimposition of forces given by the surroundings.
+Besides that, as seen in Table II, the approach does not scale in terms of performance (convergence, safety, and proximity maintenance) when the number of robots grows, due to the complex superimposition of forces given by the surroundings.
 PACNav [7] Our ($\epsilon=2.00$) Our ($\epsilon=1.052$)
-Scenario avg. time (s) SR avg. time (s) SR avg. time (s) SR
-1 55.01 1.00 39.96 1.00 24.23 1.00
-2 - 0.00 45.71 1.00 29.53 1.00
-3 - 0.00 120.33 1.00 55.87 1.00
-Table II: The simulation results obtained in terms of time to reach the goal and success rate for $10$ instances in $3$ different scenarios, comparing PACNav [7] with our method in two different parameter configurations.
+Scenario N avg. time (s) SR avg. time (s) SR avg. time (s) SR
+1 3 55.01 1.00 39.96 1.00 24.23 1.00
+2 9 - 0.00 45.71 1.00 29.53 1.00
+3 9 - 0.00 120.33 1.00 55.87 1.00
+Table II: The simulation results obtained in terms of time to reach the goal and success rate for $10$ instances in 33 different scenarios, comparing PACNav [7] with our method in two different parameter configurations.
+For further simulation results, please refer to the attached multimedia material, where we include simulations with up to $16$ UAVs in crowded conditions.
+Notice that even if the sufficient condition in Proposition 1 is not respected, the robots can still safely reach the goal in environments with $56\%$ of free space, as shown in multimedia material.
 V Experimental Results
 To validate our algorithm, we implemented it on real robotic platforms.
-We tested the proposed approach in two different scenarios: in the field with artificial obstacles, and in the forest.
-While in the former case we designed the distances between the obstacles and we relied on GNSS signal, in the forest we faced more challenging conditions both in terms of obstacle density, and in terms of localization, since we did not rely on GNSS signal.
-We started in a more controlled environment in order to have easier access to a ground truth measurement of the positions of the robots to evaluate the algorithm, and also for a higher repeatability of the experiment, because of the reduced complexity of the obstacle distribution.
-Obstacle detection
 To detect the surrounding obstacles we relied on the RPLiDAR A3.
@@
 By doing that, we exploit both the reactiveness of the raw data coming from the LiDAR pointcloud, and at the same time, the scan-matching based obstacle representation, which is provided by the Hector mapping algorithm.
-Relative localization
+At the same time we do not need to know the obstacles’ size, since we are considering the closest point, we can account for a small value for $\delta^{o}$.
 The relative positions of the neighboring robots are obtained by relying on the UltraViolet Direction And Ranging (UVDAR) technology [28].
-The sensor system operates by transmission of modulated ultraviolet light from active on-board markers.
-These markers consist of blinking ultraviolet LEDs and their signal is retrieved by on-board cameras with tailored filters.
-Each of these markers is individually identifiable according to its blinking signal, and retrieving a set of them corresponding to a specific Unmanned Aerial Vehicle (UAV) allows the system to estimate the UAV’s relative pose.
-V-A Experiments in a field
-We tested our approach on $N=3$ aerial robots in the field with artificial obstacles.
-We considered $\delta_{i}=1$ (m), $\delta^{o}=0.05$ (m), $r_{s,i}=6$ (m), $d_{1}=d_{3}=1$ (m), $d_{2}=d_{4}=1$ (m), $\beta_{i}^{D}=2.0$, $\Gamma_{ij}=12$ (m), $d_{u}=1.2$ (m), $\epsilon_{p}=\epsilon_{o}=2$, and the adjacency matrix that defines the flocking constraints is $A_{d}=\left[\begin{smallmatrix}0&0&1\\ 0&0&1\\ 1&1&0\end{smallmatrix}\right]$.
-In this first experiment, we positioned $N_{o}=7$ artificial obstacles that mimic a portion of the forest.
-The robots relied on GNSS.
-In accordance with theoretical findings, despite the presence of uncertainties in the perception, and tracking errors due to model mismatch and external disturbances, each robot safely converged to the goal regions maintaining the desired flocking constraints.
-In Figure 6, we depict five snapshots at different time instants during the mission.
-The robots are depicted with colored circles depending on the time step, in transparency we depict $\delta_{i}+d_{u}$, which represents the amount of uncertainties that we are considering.
-The obstacles are depicted in black.
-The green crosses indicate the goal positions.
-As seen, the robots respected the flocking and the safety constraints and converged successfully to the goal regions.
-Figure 6: Experimental result with $N=3$ aerial robots.
-We depicted $5$ snapshots in different time instants.
-Green crosses indicate the goal positions and the black circles indicate the obstacles.
-Transparent areas have radius equal to $d_{u}+\delta_{i}$.
-Figure 7: On-board relative localization of UAV 1 and UAV 3 from the viewpoint of the UAV 2.
-Figure 8: Distances measured by UVDAR system on UAV 2 from UAV 1 (blue) and UAV 3 (red). The flocking and safety constraints are denoted in magenta.
-In Figures 7 and 8, we show the positions of the neighboring robots from the perspective of UAV $3$.
-The data provided by the onboard sensors is quite noisy, especially when distances between robots are significant and/or when occlusions due to the obstacles or other robots occur.
-Nevertheless, from Figure 8, it is clear that safety constraints are respected.
-On the other hand, the flocking constraint appears to exceed the limit of $\Gamma_{ij}=12$ (m) at certain times.
-By plotting the distances from the odometry data (Figure 9), it is clear that these are actually jumps in the measurements and not in the actual relative positions.
-Figure 9: Distances between UAVs using odometry.
-When this situation occurs, the uncertainty of the measure can be quite high, and since we set $d_{u}=1.2$ (m) as a constant value it may cause an appearance of constraint violation.
-From a practical point of view, the neighbor position is set to be at a maximum distance of $\Gamma_{ij}-d_{u}$, in order to maintain a well-defined cell, or we simply discard that value if the associated uncertainty is unreasonably large.
-Another solution would be to increase the value of $d_{u}$, but this would be at the cost of a drop of the performance.
-Similar evaluation was done for UAV $3$ (see Figure 1011).
-In this case the flocking constraints are active on both neighbors and, as it seen in Figure 9, the constraints are satisfied.
-Figure 10: On-board relative localization of UAV 1 and UAV 2 from UAV 3 viewpoint.
-Figure 11: Distances of UAV 1 (red) and UAV 2 (blue) measured by UVDAR system on UAV 3. The flocking and safety constraints are denoted in magenta.
-V-B Experiments in a dense environment
-The second experimental deployment is conducted in the forest, where we do not have access to the GNSS signal, nor we can set a priori the obstacle positions.
-In this case, we change parameters to deal with more crowded conditions, according to the theoretical findings.
-The goal positions are set to be $40$ (m) ahead with respect the initial position of each robot, notice that we do not need to have a common reference frame, however we need to satisfy the requirement stated in Corollary 1, i.e., $\Gamma_{ij}>\|e_{i}-e_{j}\|>\Delta_{ij}$, $\forall i,j=1\dots N$ with $i\neq j$.
-We tested the algorithm with up to $N=4$, by selecting $\delta_{i}=0.3$ (m), $\delta^{o}=0.05$ (m), $r_{s,i}=6$ (m), $d_{1}=d_{3}=0.5$ (m), $d_{2}=d_{4}=1$ (m), $\beta_{i}^{D}=2.5$, $\Gamma_{ij}=12$ (m), $d_{u}=0.8$ (m), $\epsilon_{p}=\epsilon_{o}=1.052$, and the adjacency matrix that defines the flocking constraints is
-$A_{d}=\begin{bmatrix}0&1&0&1\\ 1&0&1&0\\ 0&1&0&1\\ 1&0&1&0\end{bmatrix}.$
-In Figure 12, we show the trajectory of UAV $2$, the positions of the other robots measured by UVDAR system, and the map built by UAV $2$ by using Hector SLAM during the mission.
+The sensor system operates by detection of blinking ultraviolet LEDs.
+The blinking signal is retrieved by on-board cameras with tailored filters.
+Each UAV is individually identifiable according to its blinking frequency.
+The pose of the UAVs is inferred through computer vision techniques.
+The experimental deployment is conducted in the forest, where we do not have access to the GNSS signal.
+The goal positions are set to be $40$ (m) ahead with respect to the initial position of each robot, notice that we do not need to have a common reference frame.
+However, we need to satisfy the requirement stated in Proposition 1, i.e., $\Gamma_{ij}>\|e_{i}-e_{j}\|>\Delta_{ij}$, $\forall i,j=1\dots N$ with $i\neq j$.
+We tested the algorithm with up to $N=4$, by selecting $\delta_{i}=0.3$ (m), $\delta^{o}=0.05$ (m), $r_{s,i}=6$ (m), $d_{1}=d_{3}=0.5$ (m), $d_{2}=d_{4}=1$ (m), $\beta_{i}^{D}=2.5$, $\Gamma_{ij}=12$ (m), $d_{u}=0.8$ (m), $\epsilon_{p}=\epsilon_{o}=1.052$, and the adjacency matrix that defines the flocking constraints is $A_{d}=\left[\begin{smallmatrix}0&1&0&1\\ 1&0&1&0\\ 0&1&0&1\\ 1&0&1&0\end{smallmatrix}\right]$.
+In Figure 8-(a) we show the trajectory of UAV $2$, the positions of the other robots measured by UVDAR system, and the map built by UAV $2$ by using Hector SLAM during the mission.
 In Figure 6, we depict the distances between UAV $2$ and the other robots according to the UVDAR measurements.
-Similar considerations as in the previous experiments apply also in this case.
 In Table III, we report the average velocities, the maximum velocities, and the time to reach the goal for each UAV during the mission.
-To verify the success of the mission, we report in Figure 15 the actual trajectories for all the 4 robots in the mission, while in Figure 14 we show the distances between the UAVs during the mission.
-As we expected the safety constrained and the flocking constraints (according to the adjancency matrix $A_{d}$) are satisfied.
+To verify the success of the mission, we report in Figure 8-(b) the actual trajectories for all the 4 robots in the mission, while in Figure 7 we show the distances between the UAVs during the mission.
+As we expected the safety constraints and the flocking constraints (according to the adjancency matrix $A_{d}$) are satisfied.
 To show reliability and robustness to parameter changes, we report in Table IV the results obtained from other four experiments in the forest, with $N=3$ UAVs.
@@
 Nevertheless, through the table we want to show the repeatability of the experiments and the robustness to parameter changes.
-The parameters for each experiment were set as follows: ##\##experiment $=[1,2,3,4]$, $N=[3,3,3,3]$, $\delta_{i}=[0.3,0.3,0.3,0.3]$ (m), $\delta^{o}=[0.05,0.05,0.05,0.05]$ (m) $r_{s,i}=[6,6,6,6]$ (m), $d_{1}=d_{3}=[0.5,0.5,0.5,0.5]$ (m), $d_{2}=d_{4}=[1,1,1,1]$ (m), $\beta_{i}^{D}=[2.0,2.0,1.0,0.6]$, $\Gamma_{ij}=[12,10,10,10]$ (m), $d_{u}=[0.75,0.75,0.75,0.75]$ (m), $\epsilon_{p}=\epsilon_{o}=[1.052,1.052,1.010,1.052]$, and the adjacency matrix that defines the flocking constraints is the same for all the four experiments, $A_{d}=\left[\begin{smallmatrix}0&0&1\\ 0&0&1\\ 1&1&0\end{smallmatrix}\right].$
+The parameters for each experiment were set as follows: #\#experiment =[1,2,3,4]=[1,2,3,4], $N=[3,3,3,3]$, $\delta_{i}=[0.3,0.3,0.3,0.3]$ (m), $\delta^{o}=[0.05,0.05,0.05,0.05]$ (m) $r_{s,i}=[6,6,6,6]$ (m), $d_{1}=d_{3}=[0.5,0.5,0.5,0.5]$ (m), $d_{2}=d_{4}=[1,1,1,1]$ (m), $\beta_{i}^{D}=[2.0,2.0,1.0,0.6]$, $\Gamma_{ij}=[12,10,10,10]$ (m), $d_{u}=[0.75,0.75,0.75,0.75]$ (m), $\epsilon_{p}=\epsilon_{o}=[1.052,1.052,1.010,1.052]$, and the adjacency matrix that defines the flocking constraints is the same for all the four experiments, $A_{d}=\left[\begin{smallmatrix}0&0&1\\ 0&0&1\\ 1&1&0\end{smallmatrix}\right].$
 The videos of the experiments discussed in this section can be found in the multimedia material.
@@
 Table III: Average speed, maximum speed, and time to reach the goal of each UAV in the experiment from Figure 8.
+Figure 6: Distances measured by UVDAR system on UAV 2 from UAV 1 (blue), UAV 3 (orange), and UAV 4 (red). The flocking and safety constraints are in magenta.
+Figure 7: Distances between UAVs computed based on odometries transformed to a common reference frame obtained through alignment of maps of individual UAVs in a post-prossessing phase.
+Refer to caption Refer to caption
+(a) (b)
 Figure 8: Experiment in the forest.
-The trajectory of UAV $2$ is depicted in green.
-The positions measured through UVDAR by UAV $2$ are depicted in orange (UAV $1$), blue (UAV $3$), and red (UAV $4$).
-The map is built by UAV $1$ by means of Hector mapping is also represented, in particular the free space is white, obstacles are black and the unexplored area is grey.
-Figure 13: Distances measured by UVDAR system on UAV 2 from UAV 1 (blue), UAV 3 (orange), and UAV 4 (red). The flocking and safety constraints are in magenta .
-Figure 14: Distances between UAVs using odometry.
-Figure 15: Experiment in the forest.
-The trajectory of UAV $1$ is depicted in orange, UAV $2$ in green, UAV $3$ in blue, and UAV $4$ in red.
+In (a) the trajectory of UAV $2$ is depicted in green.
+The positions measured through UVDAR by UAV 22 are depicted in orange (UAV 11), blue (UAV 33), and red (UAV 44).
 The map built by UAV $2$ by means of Hector mapping is also represented, in particular the free space is white, obstacles are black and the unexplored area is grey.
+In (b) we depicted the actual trajectories obtained in post-process phase through transformation of odometries to a common reference frame.
+The map was built by UAV 11.
 UAV1 UAV2 UAV3
@@
 VI Conclusions
-In this paper, we presented a novel distributed algorithm to swarm reliably and effectively in dense cluttered environments such as a forest, accounting for tracking errors and measurements uncertainties.
+In this paper, we presented a novel distributed algorithm to swarm reliably and effectively in dense cluttered environments represented by a forest, accounting for tracking errors and measurements uncertainties.
 We provide sufficient conditions for the theoretical guarantees of safety, proximity maintenance and convergence, compare it with state of the art algorithms, and finally we validate our solution through experiments in the field and in dense environment such as a forest.
-Our approach presents some imitations; the algorithm is limited to work on a plane, i.e., the reference in the altitude for each robot is constant and equal.
-To ensure the convergence towards the goal region, we assume that the obstacles are at a certain distance one each other (see Corollary 1).
-When these conditions are not met, providing to the algorithm an higher level re-planner can be beneficial to enhance performance.
-Future research directions will be focused on an effective 3D extension of the Lloyd algorithm and to increasing the agility of the swarm.
+Our approach presents some limitations; the algorithm is limited to work on a plane, i.e., the reference in the altitude for each robot is constant and equal.
+To ensure the convergence towards the goal region, we assume that the individual obstacles are separated by a certain minimum distance (see Proposition 1, notice that these are just sufficient conditions and not necessary).
+Providing to the algorithm a higher level re-planner can be beneficial to enhance performance.
+Future research directions will be focused on an effective 3D extension of the Lloyd algorithm and to increase the agility of the swarm.
 CNN

### v3

@@
-Distributed Lloyd-Based Algorithm for Uncertainty-Aware Multi-Robot Under-Canopy Flocking
+arXiv:2504.18840v3[cs.RO] 27 Feb 2026
 Abstract
 In this letter, we present a distributed algorithm for flocking in complex environments that operates at constant altitude, without explicit communication, no a priori information about the environment, and by using only on-board sensing and computation capabilities.
-We provide sufficient conditions to guarantee that each robot reaches its goal region in a finite time, avoiding collisions with obstacles and other robots without exceeding a desired maximum distance from a predefined set of neighbors (flocking or proximity constraint).
-The proposed approach allows to operate in crowded scenarios and to deal with tracking errors and on-board sensing errors, without violating safety and proximity constraints.
+We provide sufficient conditions to guarantee collision avoidance with obstacles and other robots without exceeding a desired maximum distance from a predefined set of neighbors (flocking or proximity maintenance constraint) during the mission.
+The proposed approach allows to operate in crowded scenarios and to explicitly deal with tracking errors and on-board sensing errors.
 The algorithm was verified through simulations with varying number of UAVs and also through numerous real-world experiments in a dense forest involving up to four UAVs.
@@
 Multi-robot systems, Distributed control, Lloyd-based algorithms.
-Video—https://mrs.fel.cvut.cz/rbl-wild-2024
-Github—https://github.com/ctu-mrs/rbl_controller
+Video—https://mrs.fel.cvut.cz/rbl
+Github—https://github.com/ctu-mrs/rbl
 I Introduction
@@
 For many of these applications, having a swarm of robots able to flock together towards a desired location can be highly beneficial, due to redundancies and parallelization.
-The definition of flocking is provided by the seminal work of Reynolds [1], which defines it as a non-colliding aggregate motion.
 Depending on the environments where robots operate, the reliability or availability of explicit communication between robots may be compromised.
-In this work, we propose a Lloyd-based algorithm that relies only on on-board sensors.
-We aim to synthesize a distributed algorithm, where each robot safely reaches the desired position, while maintaining proximity between robots in the team.
-We provide sufficient conditions to guarantee safety, proximity maintenance and convergence towards the goal in cluttered environments by accounting for bounded tracking errors and uncertainties in the estimation of neighbors and obstacles’ positions.
+In this work, we propose a Lloyd-based algorithm that relies solely on on-board sensing.
+We have synthesized a distributed algorithm in which each robot safely reaches its desired position while maintaining proximity to its teammates in cluttered environments.
+We choose Lloyd algorithm as a base for our method as it enables seamless multi-robot coordination without requiring robots to share unobservable or global information, and its safety guarantees are not dependent on finely tuned parameters.
 Our algorithm takes inspiration from the nature; just like flock of birds or a school of fish, the robots do not need a leader and each robot makes its own decisions based on relative information, without relying on explicit communication.
@@
 Among them, a smaller subset is devoted to applications in the field without relying on explicit communication between the robots.
-To position our algorithm within the state of the art, we consider the following features.
+To position our work within the state of the art, we consider the following features.
 Distributed control: each robot makes decisions on the basis of local information, which is obtained by relying only on on-board sensors.
@@
 Safety: collision and obstacle avoidance is guaranteed at any time.
-Convergence: the sufficient conditions for the robots to converge to their goal locations are provided.
 Proximity maintenance: the distance between robots is guaranteed to be below the set threshold on maximum distance.
-Uncertainties: the system is provably safe and respects proximity constraints even in the presence of bounded tracking errors and/or bounded measurement uncertainties.
+Uncertainties: the approach is robust to bounded tracking errors and bounded measurement uncertainties.
 Cluttered and highly dense environments: the algorithm can operate in cluttered environments such as a forest.
-In particular, in absence of uncertainties and convergence guarantees, the robots can operate in environments where the minimum distance between two obstacles is large enough to allow the single robot to pass.
-To ensure convergence and account for uncertainties, a sufficient condition is that the obstacles are spaced far enough to allow passage of twice the robot’s size.
 On-board localization: the algorithm is verified in unknown environments relying only on on-board localization.
@@
 Safety $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\times$
-Convergence $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
 Uncertainties $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\times$ $\circ$ $\circ$ $\circ$
@@
 As shown in Table I, our algorithm is unmatched in fulfilling all the specified features of interest.
-In fact, despite [5, 4] proposing valid solutions to the problem, their main limitations are the requirement of a trajectory broadcasting network, the absence of both convergence analysis and a rigorous account of tracking errors and uncertainties.
+In fact, with respect to [9], we included the proximity maintenance constraints, the ability to deal with cluttered environments, tracking errors and uncertainties, without relying on explicit communication, nor an external localization system.
+Despite [30, 29] proposing valid solutions to the problem, their main limitations are the requirement of a trajectory broadcasting network and the absence of account of proximity maintenance, tracking errors and uncertainties.
 Similarly, in [31] the robots need to broadcast information such as poses, velocities, poses covariance, and global extrinsic transformations.
-On the other hand, [7, 8], do not rely on an explicit communication network, but they exhibit degraded performance in crowded scenarios, lack of safety, proximity maintenance, convergence guarantees analysis, and robustness to tracking errors and uncertainties.
+On the other hand, [2, 17], do not rely on an explicit communication network, but they exhibit degraded performance in crowded scenarios, lack of safety guarantees, proximity maintenance, and robustness to tracking errors and uncertainties.
 In [32], the authors present a distributed solution, which in theory does not rely on explicit communication and accounts for uncertainties.
@@
 In Table I, we report additional relevant state-of-the-art algorithms.
-Notably, all of them share limitations related to the use of an explicit communication network and/or to not rigorously account for safety, convergence, proximity maintenance and the uncertainties.
+Notably, all of them share limitations related to the use of an explicit communication network and/or to not account for safety, proximity maintenance and the uncertainties.
 Conversely, in our work, we provide a method which implements all the fore-mentioned features as well as a comprehensive rule for the selection of the algorithm’s parameters based on measurable quantities (e.g., tracking error, or covariance of state estimates), which is often an unclear or empirical process in the state-of-the-art algorithms.
@@
 The key novelties lie in our use of an adaptive control strategy to compensate for tracking errors and measurement uncertainties, combined with Convex Weighted Voronoi Diagrams (CWVD) to enhance the performance in a dense forest scenarios.
-Second, we provide sufficient conditions for safety, proximity maintenance, and convergence in cluttered environments and in the presence of flocking constraints.
-Finally, we validate the algorithm in the forest, relying only on on-board sensors, i.e., Inertial Measurement Unit (IMU), 2-D Lidar and UVDAR system [28], without relying neither on an explicit communication network nor GNSS signal.
+Second, we provide sufficient conditions for safety and proximity maintenance in cluttered environments and in the presence of flocking constraints.
+Finally, we validate the algorithm in the forest, relying only on on-board sensors, i.e., Inertial Measurement Unit (IMU), 2-D Lidar, altitude sensor, and UVDAR system [26], without relying neither on an explicit communication network nor GNSS signals nor having a common reference frame.
 II Problem description
@@
 Problem 1.
-Let $N$ be the number of robots in the swarm, let $\mathcal{N}_{i}$ be the set of neighbors of the robot $i$, and $\bar{\mathcal{N}}_{i}\subseteq\mathcal{N}_{i}$ the set of neighbors that has to remain connected the entire mission.
-We want to synthesize a distributed control law, such that the swarm maintains the desired proximity constraints, which is characterized by a set of maximum distances $\Gamma_{ij}=\Gamma_{ji},\,\,\,\forall i\in{1,\dots,N},\forall j\in\bar{\mathcal{N}}_{i}$.
+Let $N$ be the number of robots in the swarm, let $\mathcal{N}_{i}$ be the set of neighbors of the robot $i$, and $\bar{\mathcal{N}}_{i}\subseteq\mathcal{N}_{i}$ the set of neighbors that has to remain in the proximity of the robot $i$ for the entire mission.
+We want to synthesize a distributed control law, such that the swarm maintains the desired proximity constraints, which is characterized by a set of maximum distances $\Gamma=\{\Gamma_{ij}\mid i\in{1,\dots,N},j\in\bar{\mathcal{N}}_{i}\}$.
 Each robot in the swarm has to reach its goal region avoiding collisions with other robots and obstacles.
@@
 $J_{\operatorname{cov}}({p},\mathcal{V})=\sum_{i\in\mathcal{N}_{i}}\int_{\mathcal{V}_{i}}\left\|q-p_{i}\right\|^{2}\varphi_{i}(q)dq,$ (1)
-where $\mathcal{V}=\{\mathcal{V}_{1},\dots,\mathcal{V}_{N}\}$, $\mathcal{V}_{i}=\mathcal{V}_{i}^{p}\cap\mathcal{V}_{i}^{o}$ with $\mathcal{V}_{i}^{p}$ being the ii–th Convex Weighted Voronoi Diagram (CWVD) introduced in [13], which accounts only for neighboring robots’ positions
+where $\mathcal{V}=\{\mathcal{V}_{1},\dots,\mathcal{V}_{N}\}$, $\mathcal{V}_{i}=\mathcal{V}_{i}^{p}\cap\mathcal{V}_{i}^{o}$ with $\mathcal{V}_{i}^{p}$ being the ii–th Convex Weighted Voronoi Diagram (CWVD) introduced in [6], which accounts only for neighboring robots’ positions, while $\mathcal{V}_{i}^{o}$ is the ii–th CWVD accounting only for neighboring obstacles’ positions:
 $\small\begin{split}\mathcal{V}_{i}^{p}=\{q\in\mathcal{Q}\mid w_{i}^{x}\cos\alpha_{ij}+w_{i}^{y}\sin\alpha_{ij}<\frac{1}{\epsilon_{p}}&\left\|p_{i}-p_{j}\right\|,\\ &\forall j\in\mathcal{N}_{i}\}.\end{split}$ (2)
-While $\mathcal{V}_{i}^{o}$ is the ii–th CWVD accounting only for neighboring obstacles positions
 $\small\begin{split}\mathcal{V}_{i}^{o}=\{q\in\mathcal{Q}\mid w_{i}^{x}\cos\alpha_{ij}+w_{i}^{y}\sin\alpha_{ij}<\frac{1}{\epsilon_{o}}&\left\|p_{i}-o_{j}\right\|,\\ &\forall j\in\mathcal{O}_{i}\},\end{split}$ (3)
-where $w_{i}=[w_{i}^{x},w_{i}^{y}]^{\top}=q-p_{i}$ and $\alpha_{ij}=\arctan\left(\frac{y_{j}-y_{i}}{x_{j}-x_{i}}\right)$, while $\epsilon_{p(o)}\in[1,2]$, is a parameter, which regulates the cautiousness in the robot motion (i.e., $\epsilon_{p(o)}=1$ more aggressive behavior, $\epsilon_{p(o)}=2$ more conservative behavior, notice that the latter coincides with the classic Voronoi diagram).
-Figure 2: Representation of the Convex Weighted Voronoi Diagram (CWVD) $\mathcal{V}_{i}^{p}$ in (2) with $\epsilon_{p}=2$ and $\epsilon_{p}=1$.
-Where the green circle is the i–th robot and the cyan circles represents the neighboring robots.
-In Figure 2, we depict an example of the CWVD (2) for the case where $\epsilon_{p}=2$ and $\epsilon_{p}=1$.
+where $w_{i}=[w_{i}^{x},w_{i}^{y}]^{\top}=q-p_{i}$ and $\alpha_{ij}=\arctan\left(\frac{y_{j}-y_{i}}{x_{j}-x_{i}}\right)$, while $\epsilon_{p(o)}\in[1,2]$ regulates the cautiousness in the robot motion (i.e., $\epsilon_{p(o)}=1$ more aggressive behavior, $\epsilon_{p(o)}=2$ more conservative behavior, notice that the latter coincides with the classic Voronoi diagram).
+Refer to caption Refer to caption
+(a) (b)
+Figure 2: (a) Representation of the Convex Weighted Voronoi Diagram (CWVD) $\mathcal{V}_{i}^{p}$ in (2) with $\epsilon_{p}=2$ and $\epsilon_{p}=1$.
+Where the green circle is the i–th robot and the cyan circles represent the neighboring robots.
+(b) Example of $\mathcal{F}_{i}$ and $\mathcal{A}_{i}$ cell geometry with $\epsilon_{p}=\epsilon_{o}=2$.
+Red solid lines indicate the proximity maintenance constraints.
+Obstacles are represented as grey circles, while the robots are depicted in cyan.
+In Figure 2a, we depict an example of the CWVD (2) for the case where $\epsilon_{p}=2$ and $\epsilon_{p}=1$.
 By following the gradient descent of the cost function (1), we obtain the following proportional control law
@@
 Theorem 1 (Safety).
-By imposing the control $\dot{p}_{i}(\mathcal{A}_{i})$ in (4), with $\tilde{\mathcal{V}}_{i}^{p}\cap\tilde{\mathcal{V}}_{i}^{o}\cap\mathcal{S}_{i}$, collision avoidance is guaranteed at every instant of time.
+By imposing the control $\dot{p}_{i}(\mathcal{A}_{i})$ in (4), with $\mathcal{A}_{i}=\tilde{\mathcal{V}}_{i}^{p}\cap\tilde{\mathcal{V}}_{i}^{o}\cap\mathcal{S}_{i}$, collision avoidance is guaranteed at every instant of time.
 Proof.
@@
 Due to the fact that the set $\mathcal{A}_{i}$ is always convex, and by definition $p_{i},c_{\mathcal{A}_{i}}\in\mathcal{A}_{i}$, it follows that since the straight path from the robot position to the centroid position is a safe path, then applying $\dot{p}_{i}(\mathcal{A}_{i})$ in (4) preserves safety.
+Due to the robots’ size, the condition that the straight path from the robot to its centroid position is a safe path can be violated only if the $i$-th robot is at a distance greater than $2\Delta_{ij}$ ($2\Delta^{o}_{ij}$) from the $j$-th robot ($j$-th obstacle), according to (6), (7).
+Nevertheless, to experience collision the distance between the $i$-th robot and the $j$-th robot ($j$-th obstacle) needs to be less than $2\Delta_{ij}$ ($2\Delta^{o}_{ij})$, i.e., equal to $\Delta_{ij}$ ($\Delta^{o}_{ij}$), hence safety is preserved.
 ∎
-We further apply an additional reshaping of the cell geometry in order to maintain the desired proximity of the agents, i.e., to achieve swarming behavior.
+We further apply an additional reshaping of the cell geometry in order to maintain the desired proximity of the agents, i.e., to achieve flocking behavior.
 By imposing our final cell geometry $\mathcal{F}_{i}=\mathcal{A}_{i}\cap\mathcal{M}_{i}$, where
 $\mathcal{M}_{i}=\{q\in\mathcal{Q}\mid\|q-p_{j}\|\leq\Gamma_{ij},\forall j\in\bar{\mathcal{N}}_{i}\},$ (9)
-we obtain proximity maintenance while preserving safety. In Figure 3, we depict an example of cell geometry $\mathcal{F}_{i}$.
-Figure 3: Example of $\mathcal{F}_{i}$ and $\mathcal{A}_{i}$ cell geometry with $\epsilon_{p}=\epsilon_{o}=2$.
-Dashed lines indicate the proximity constraints.
-Obstacles are represented as grey circles, while the robots are depicted in cyan.
-Theorem 2 (Proximity maintenance).
+we obtain proximity maintenance while preserving safety. In Figure 2b, we depict an example of cell geometry $\mathcal{F}_{i}$.
+Theorem 2 (Proximity maintenance or flocking constraint).
 By imposing the control $\dot{p}_{i}(\mathcal{F}_{i})$ in (4), both safety and proximity maintenance are guaranteed at every instant of time.
@@
 III-B The weighting function $\varphi_{i}(q)$
-Although ensuring safety and proximity maintenance only involves properly shaping the cell geometry, achieving convergence and enhancing navigation performance largely hinges on how the function $\varphi_{i}(q)$ is defined.
-We adopt a Laplacian-shaped function centered at $\bar{p}_{i}$, which represents a point in the mission space that ultimately needs to converge to the goal location.
+Although ensuring safety and proximity maintenance only involves properly shaping the cell geometry, convergence to the goal and enhancing navigation performance largely hinges on how the function $\varphi_{i}(q)$ is defined.
+As originally proposed in [9], we adopt a Laplacian-shaped function centered at $\bar{p}_{i}$, which represents a point in the mission space that ultimately needs to converge to the goal location.
 We design the $\varphi_{i}(q)$ function as follows:
 $\varphi_{i}(q)=\text{exp}{\left(-\frac{\|q-\bar{p}_{i}\|}{\beta_{i}}\right)},$ (10)
-where the spreading factor $\beta_{i}\in\mathbb{R}$ is time-variant, it follows the dynamics
-$\dot{\beta}_{i}(\mathcal{A}_{i})=\begin{cases}-\beta_{i}\,\,\,&\text{if}\,\,\|c_{\mathcal{A}_{i}}-p_{i}\|<d_{1}\,\land\\ &\|c_{\mathcal{A}_{i}}-c_{\mathcal{S}_{i}}\|>d_{2}\\ -(\beta_{i}-\beta^{D}_{i})&\text{otherwise,}\end{cases}$ (11)
+where the spreading factor $\beta_{i}\in\mathbb{R}$ is time-variant. It follows the dynamics
+$\dot{\beta}_{i}(\mathcal{A}_{i})=\begin{cases}-k_{\beta}\beta_{i}\,\,\,&\text{if}\,\,\|c_{\mathcal{A}_{i}}-p_{i}\|<d_{1}\,\land\\ &\|c_{\mathcal{A}_{i}}-c_{\mathcal{S}_{i}}\|>d_{2},\\ -k_{\beta}(\beta_{i}-\beta^{D}_{i})&\text{otherwise,}\end{cases}$ (11)
 and
 […]
-The parameters $\beta_{i}^{D},d_{1},d_{2},d_{3},d_{4}$ are positive scalars, $R(\theta)$ is the rotation matrix, $e_{i}$ is the goal position for the ii–th robot, $\varepsilon$ is a small number, which allows to avoid rotations equal to $\pi/2$, while $\bar{{c}}_{\mathcal{A}_{i}}$ is the centroid position computed over the cell $\mathcal{A}_{i}$ with $\bar{p}_{i}\equiv e_{i}$.
-Proposition 1 (Convergence in cluttered environment and with flocking constraints).
-Let us assume we have an unbounded space cluttered with circular obstacles of radius $\delta^{o}<\Gamma_{ij}/2-2\delta$, where $\delta=\delta_{i},\forall i=1,\dots,N$, and at a minimum distance of $\|o_{i}-o_{j}\|>4\delta+2\delta^{o}$ one to the other.
-The initial and goal positions are adequately apart from each other, i.e., $\Gamma_{ij}>\|p_{i}(0)-p_{j}(0)\|>\Delta_{ij}$, $\Gamma_{ij}\geq\|e_{i}-e_{j}\|>\Delta_{ij}$, $\forall i,j=1\dots N$ with $i\neq j$.
-By imposing the control law $\dot{p}_{i}(\mathcal{F}_{i})$ in (4), (10), (11), and (12) to compute the centroid position in (5), by properly selecting $d_{1},d_{2},d_{3},d_{4}$, and $\beta^{D}_{i}$, we have that $\lim_{t\rightarrow\infty}\|p_{i}(t)-e_{i}\|<r_{s,i},\forall i=1,\dots,N.$
-Figure 4: The figure shows two possible deadlock conditions: case 1.
-is due to the small inter-space between obstacles.
-case 2.
-is due to the too short $\Gamma_{ij}$ value (or too big $\delta^{o}$ value).
-We depict in cyan the robots and in grey the static obstacles.
-The crosses represent the final goal positions while the dashed lines represent the proximity constraint topology.
-Proof.
-The proof provided in [2] does not consider presence of static obstacles nor proximity constraints.
-By including both of them, we recognize, beside the trivial initial and final conditions requirements, the addition of two potential deadlock scenarios, which are depicted in Figure 4.
-Notice that the addition of static obstacles is equivalent to non-cooperative static robots.
-In order to not affect the proof of convergence obtained in [2], two requirements need to be considered.
-First, we require a minimum distance between the obstacles of $\|o_{i}-o_{j}\|>4\delta+2\delta^{o},\,\forall i\neq j$, in order to avoid the presence of narrow hallways that can generate deadlock issues (left hand side of Figure 4).
-This is a consequence of the results in [2].
-In fact, static obstacles can be thought as robots that reached their goal and have a value of $\beta_{i}\rightarrow 0$.
-However, from [2], by design this is not possible if the robots are at a distance from their goal smaller than $d_{2}$.
-Since $d_{2}>2\delta$ and $\|e_{i}-e_{j}\|>\Delta_{ij}$, then the condition $\|o_{i}-o_{j}\|>4\delta+2\delta^{o},\,\forall i\neq j$ is equivalent.
-Second, we require that the maximum radius of the obstacles is $\delta^{o}<\Gamma_{ij}/2-2\delta$, in order to avoid strict proximity maintenance constraints (right hand side of Figure 4).
-By imposing these additional conditions, the guarantees of convergence in [2] holds.
-∎
+The parameters $\beta_{i}^{D},d_{1},d_{2},d_{3},d_{4},k_{\beta},k_{e}$ are positive scalars, $R(\theta)$ is the rotation matrix, $e_{i}$ is the goal position for the ii–th robot, $\varepsilon$ is a small number, which allows to avoid rotations equal to $\pi/2$, while $\bar{{c}}_{\mathcal{A}_{i}}$ is the centroid position computed over the cell $\mathcal{A}_{i}$ with $\bar{p}_{i}\equiv e_{i}$.
+Equations (11) and (12) specify two adaptive rules.
+Intuitively, (11) regulates the greediness of the robot $i$ in reaching its goal region, while (12) introduces a right (or left) hand-side rule to enforce a common behavioral convention across the robots and hence enhance inter-robot coordination.
 III-C Dealing with dynamics and uncertainties
-The real robot deployment requires to consider a set of additional issues.
+The robot deployment requires to consider a set of additional issues.
 First, more complex dynamics of the robots must be applied as $\dot{p}_{i}=u_{i}$ is not directly applicable.
-Second, tracking error, which can be caused by external disturbances or model mismatches, generating a difference between the desired and the actual velocity, has to be taken in to account.
+As a consequence, tracking error, which is caused by model mismatch and external disturbances, generating a difference between the desired and the actual velocity, has to be taken into account.
 Finally, we have to consider measurement uncertainties arising from limitations of the perception system providing an estimate of the positions of other robots and obstacles with a certain error.
-To account for the dynamic constraints of the robots, we synthesize the control inputs through an MPC formulation (similarly to [10, 2]), which steers each ii–th robot towards the centroid position $c_{\mathcal{F}_{i}}$, by constraining the trajectory to not exceed its $\mathcal{F}_{i}$ cell.
-By doing that, we can account for the robot dynamics without violating the safety and proximity maintenance constraints.
+To account for the dynamic constraints of the robots, we synthesize the control inputs through an MPC formulation (similarly to [32, 9]).
 Hence, each robot $i$ solves the following optimization problem (for the sake of clarity we discard the $i$ indexing),
-$\operatorname*{\min\text{imize}}_{{x}_{k},{u}_{k}}$ $\sum^{N_{t}}_{k=0}J({x}_{k},{u}_{k},c_{\mathcal{F}}),$ (13a)
+$\operatorname*{\text{minimize}}_{{x}_{k},{u}_{k}}$ $\sum^{N_{t}}_{k=0}J({x}_{k},{u}_{k},c_{\mathcal{F}}),$ (13a)
 s.t.: ${x}_{k}=f({x}_{k-1},{u}_{k-1}),\,\,\,\forall k\in(1,N_{t}),$ (13b)
 $\,{x}_{k}\in\mathcal{X},\,{u}_{k}\in\mathcal{U},\,\,\,\forall k\in(0,N_{t}),$ (13c)
-$p_{k}\in\mathcal{F},\,\,\,\forall k\in(0,N_{t}),$ (13d)
-${x}_{0}={x}_{\textrm{init}},$ (13e)
-where $k$ indicates the time index, $N_{t}$ indicates the horizon length, $J$ represents the cost function, ${x}_{k}$ and ${u}_{k}$ are the state and control inputs respectively, $f({x}_{k-1},{u}_{k-1})$ indicates the dynamics of the robot, ${x}_{k}\in\mathcal{X},\,{u}_{k}\in\mathcal{U}$ are the state and input constraints respectively, $p_{k}\in\mathcal{F}$ constrains the future positions of the robot in its cell, and ${x}_{0}={x}_{\textrm{init}}$ is the initial condition.
+${x}_{0}={x}_{\textrm{init}},$ (13d)
+where $k$ indicates the time index, $N_{t}$ indicates the horizon length, $J$ represents the cost function, ${x}_{k}=[p_{k}^{\top},\dot{p}_{k}^{\top},\ddot{p}_{k}^{\top}]^{\top}$ and ${u}_{k}$ are the state and control inputs respectively, $f({x}_{k-1},{u}_{k-1})$ indicates the dynamics of the robot, ${x}_{k}\in\mathcal{X},\,{u}_{k}\in\mathcal{U}$ are the state and input constraints respectively, and ${x}_{0}={x}_{\textrm{init}}$ is the initial condition.
 Both the tracking errors and measurement uncertainties can be accounted for by means of an adaptive strategy.
-In fact, it can be achieved by adapting the value of $\beta_{i}$ in (11) in an online fashion.
+Specifically, it is achieved by adapting the value of $\beta_{i}$ in (11) in an online fashion.
 An alternative approach would be reshaping the $\mathcal{F}_{i}$ cell, similarly to [32].
@@
 More formally, it can be done by imposing a lower bound to $\beta_{i}$ in (11) equal to
-$\beta^{\min}_{i}=\operatorname*{arg\,min}_{\beta_{i}}\left(\|c_{\mathcal{F}_{i}}-p_{\partial\mathcal{F}_{i}}\|-d_{u}\right)^{2},$
-where $p_{\partial\mathcal{F}_{i}}$ is the position of the closest point on the contour of $\mathcal{F}_{i}$, while $d_{u}$ is a threshold value that has to be tuned on the basis of the tracking errors and measurement uncertainties.
+$\beta^{\min}_{i}=\operatorname*{arg\,min}_{\beta_{i}}\left(\|c_{\mathcal{F}_{i}}-p_{\partial\mathcal{F}_{i}}\|-d_{u}\right)^{2},$ (14)
+where $p_{\partial\mathcal{F}_{i}}$ is the position of the closest point to $c_{\mathcal{F}_{i}}$ on the boundary of $\mathcal{F}_{i}$, while $d_{u}$ is a threshold value that has to be tuned on the basis of the tracking errors and measurement uncertainties.
 For the sake of clarity, we depicted three scenarios in Figure 3.
@@
 Notice that the dimension of the region $\mathcal{Z}_{i}$ can be controlled by $d_{u}^{t}$.
-To be conservative, the value of $d_{u}^{t}$ should be determined experimentally as the maximum tracking error, and it can be set as a function of the velocity.
+To be conservative, the value of $d_{u}^{t}$ should be determined empirically as the maximum tracking error, and it can be set as a function of the velocity.
 In Figure 3b, we show the case where we have measurement uncertainties.
@@
 In this case, $\mathcal{K}_{i}$ indicates the region where the centroid $c_{\mathcal{F}_{i}}$ has to be constrained to preserve safety.
-By relying on geometric reasoning, depending on the measurement uncertainty, when $\left(1-\frac{1}{\epsilon_{p}}\right)\|p_{i}-p_{j}\|<k\sqrt{\lambda^{\max}}+\delta$, we can tune the parameter $d_{u}^{m}$ as
-$d_{u}^{m}\geq\left\|\left(1-\frac{1}{\epsilon_{p}}\right)\|p_{i}-p_{j}\|-\left({k\sqrt{\lambda^{\max}}}+\delta\right)\right\|+\delta,$ (14)
-where $\lambda^{\max}$ is the maximum eigenvalue of the covariance matrix associated with the measurement and $k$ is a scale factor that defines the confidence ellipse (e.g., $k\approx 2.795$ for a confidence level of $95\%$).
-From a practical point of view, the value of $d_{u}^{m}$ may be set empirically to a constant that satisfies inequality (14) in the worst-case expected scenario.
+Let us consider the case where $\epsilon_{p(o)}=1$.
+By relying on geometric reasoning (refer to Figure 3b), depending on the inter-robot distance measurement uncertainty, we can select the parameter $d_{u}^{m}$ as
+$d_{u}^{m}\geq k\sqrt{\lambda^{\max}}+\delta_{j}+\delta_{i},$ (15)
+where $\lambda^{\max}$ is the maximum eigenvalue of the covariance matrix associated with the distance measurement and $k$ is a scale factor that defines the confidence ellipse.
+From a practical point of view, the value of $d_{u}^{m}$ may be set empirically to a constant that satisfies inequality (15) in the worst expected scenario.
 In real-world equipments the uncertainties associated with the measurements are far from being constant values.
@@
 This technique offsets the set-points of the robots from an ideal target location according to the probability distributions of relative pose observations.
-On the other hand, to preserve proximity maintenance, if $\Gamma_{ij}-\|p_{i}-p_{j}\|\approx k\sqrt{\lambda^{\max}}$, we need to set $d_{u}^{m}\geq k\sqrt{\lambda^{\max}}$.
-Since this value may be quite high in practice, it is sufficient to impose that if the mean value of the neighbor position is out of the proximity area, then we can re-project its position to the closest point inside the proximity area.
-In this way, the proximity to the neighbors is effectively maintained.
-Finally, we show the combination of tracking error and measurement uncertainties, where $d_{u}=d_{u}^{m}+d_{u}^{t}$ in Figure 5c.
+On the other hand, to preserve proximity maintenance, a sufficient condition is that $d_{u}^{m}\geq k\sqrt{\lambda^{\max}}$.
+Finally, the tracking error and measurement uncertainties are combined as $d_{u}=d_{u}^{m}+d_{u}^{t}$ in Figure 3c.
+Hence, by imposing the control $\dot{p}_{i}(\mathcal{F}_{i})$ in (4), with $\mathcal{F}_{i}=\tilde{\mathcal{V}}_{i}^{p}\cap\tilde{\mathcal{V}}_{i}^{o}\cap\mathcal{S}_{i}\cap\mathcal{M}_{i}$ and assuming a maximum uncertainty in the estimation of the other robots’ distances equal to $k\sqrt{\lambda^{\max}}$, by selecting $d_{u}^{m}\geq k\sqrt{\lambda^{\max}}$, and assuming a bounded maximum tracking error less or equal than $d^{t}_{u}$, both safety and proximity maintenance are preserved by selecting $\beta_{i}^{\min}$ according to (14) for all $i=1,\ldots,N$.
+Specifically, let us choose $d_{u}^{m}$ such that it satisfies (15) and $d_{u}=d^{t}_{u}+d_{u}^{m}$.
+By enforcing a lower bound $\beta_{i}^{\min}$ (14), we set a minimum distance $d_{u}$ between the centroid position $c_{\mathcal{F}_{i}}$ and the boundary of the cell $\partial\mathcal{F}_{i}$.
+Due to the definition of $d_{u}$, this margin explicitly accounts for both estimation uncertainty and tracking error, thereby ensuring safety and proximity maintenance for all admissible realizations of the neighbors’ positions within the confidence ellipse.
+More intuitively, the centroid is constrained to remain within a region that preserves safety and proximity maintenance by accounting for tracking error and all possible neighbors’ positions consistent with the uncertainty bounds.
+If there exists no point $q\in\mathcal{F}_{i}$ such that $\text{dist}(q,\partial\mathcal{F}_{i})\geq d_{u}$, according to (14), the centroid position $c_{\mathcal{F}_{i}}$, is the point that maximizes the distance from $\partial\mathcal{F}_{i}$.
+Under this condition, strict guarantees of safety and proximity maintenance cannot be ensured for all possible uncertainty realizations.
+Nevertheless, in this case the robot adopts the most conservative behavior, placing the centroid at the location that maximizes boundary clearance.
+The guarantees provided by the proposed conditions are therefore sufficient and hold whenever a feasible margin exists; they rely on worst-case realizations within the assumed uncertainty bounds and bounded tracking errors.
 Figure 3: Dealing with uncertainties: (a) tracking error case, (b) measurement uncertainty case, (c) combined tracking error and measurement uncertainty.
@@
 In practice, it does not matter how $d_{u}^{m}$ and $d_{u}^{t}$ are chosen, what matters is the sum $d_{u}$, hence in the case of underestimation of one of the two terms, it is compensated by the other.
-Remark 2 (Convergence for the adaptive solution).
-By relying on the adaptive solution, while safety and proximity maintenance have been taken into account, we lost the convergence guarantees.
-In fact, by relying on the adaptive strategy that accounts for tracking error and measurement uncertainties, a new sufficient condition is introduced, which strictly depends on $d_{u}$.
-In particular, given the minimum distance between the obstacles in the environment being $d^{\min}_{o}$, in order to ensure the passage of the robot between two obstacles, it can be shown through geometrical reasoning, that we need to have $0\leq d_{u}<\frac{d^{\min}_{o}}{2}-(\delta_{i}+\delta^{o})$.
-A situation where the cell is too small to allow the centroid to be at the required distance from the boundary of the cell can also occur.
-However, we claim that this is not an issue for safety nor for proximity maintenance, since in these cases the robot velocity approaches zero.
-Such situation is transitory and may occur in very crowded conditions or when the robot has to maintain proximity with two or more robots that are close to the maximum allowed distance $\Gamma_{ij}$.
 IV Simulation results
-In this section we show through simulations, the scalability, repeatability and robustness of our approach, and we provide a comparison with a state-of-the-art algorithm.
-In particular, we report a comparison with PACNav [7], an algorithm that does not rely on communication network, and shares similar features with the proposed algorithm (refer to Table I).
+In this section, we first show through simulations the scalability, repeatability and robustness of our approach in two different scenarios, with and without accounting for uncertainties and tracking errors.
+Then we report a comparison with PACNav [2], the algorithm does not rely on communication network, and shares similar features with the proposed algorithm (refer to Table I).
 Notice that comparing with algorithm such as [30, 29, 31] is out of scope of this paper.
-Our method does not perform better than algorithms that can share information, such as future trajectories, among the robots.
+In most of the scenarios, our method does not perform better than algorithms that can share information, such as future trajectories, among the robots.
 Despite [2] performs well in relatively simple environments and with a limited number of robots, it lacks reliability in more complex scenarios.
 In addition, crucial requirements such as safety and proximity maintenance depend strongly on fine parameter tuning.
-This is a common issue in the state-of-the-art algorithms, which is overcome in our solution, where the parameter selection affects only the performance of the mission, while, safety, proximity maintenance and convergence are met if they are selected according to the theoretical findings.
-We perform the comparison in the MRS simulator [32] in the same environment and same model dynamics for both algorithms.
-We simulate three scenarios with different number of robots and with an increasingly complex obstacle configurations.
+This is a common issue in the state-of-the-art algorithms, which is overcome in our solution, where the parameter selection affects only the performance of the mission, while, safety and proximity maintenance are met if they are selected according to the theoretical findings.
+All the simulations are performed by relying on the MRS simulator [3], which account for full rigid-body UAV dynamics.
+In the first simulations, we consider different set of parameters $d_{u}=[0,1.0]$ (m) in two different scenarios, with $N=[9,16]$, $\delta_{i}=0.2$ (m), $\epsilon=1.0$, and randomly generated circular obstacles with $\delta_{o}=0.15$ (m) (see Figure 4).
+We select the parameters as follows for all the simulations $r_{s,i}=5.0$ (m), $d_{1}=d_{2}=d_{3}=d_{4}=1$ (m), $\beta_{i}^{D}=0.15$, $\Gamma_{ij}=10$ (m), while the adjacency matrices defining proximity maintenance constraints are defined as grid matrices, as depicted in Figure 4 (red lines).
+We simulated UVDAR noise as a bounded radial random measurement error with magnitude not exceeding $0.8$ (m).
+In Table II, we report the quantitative results in terms of average time to reach the goal and success rate over a total of $40$ runs.
+As expected, setting $d_{u}=0$ (m) leads to better performance in terms of time to reach the goal; however, it does not ensure collision avoidance between the robots.
+In contrast, selecting $d_{u}$ according to the theoretical findings results in safe behavior and preserves the proximity maintenance constraints.
+Figure 4 illustrates the two scenarios with $N=[9,16]$ robots, and shows the trajectories generated for one instance with $d_{u}=1.0$ (m).
+Figure 5 illustrates the distance between each robot in the two scenarios.
+The colored lines indicate whether the proximity maintenance constraint is active.
+To further validate the robustness of the proposed approach, we additionally performed 200 simulations across four different parameter sets in the same scenario depicted in Figure 4a, consistently achieving the goal preserving safety and the proximity maintenance constraints.
+Refer to caption Refer to caption
+(a) (b)
+Figure 4: Simulation results with $N=9$ and $N=16$ robots, with $d_{u}=1.0$ (m).
+The images represent example scenarios used for generation of results presented in Table II.
+Colored circles represent the robots’ positions at different time $t$.
+Red lines represent the proximity maintenance constraints between the robots.
+Black dots are the obstacles in the scene.
+Figure 5: Distances between UAVs in the simulations in Figure 4a and 4b.
+The flocking and safety constraints are in magenta.
+Colored lines indicate the presence of an active flocking constraint between the two robots.
+$d_{u}=0\penalty 10000\ (m)$ $d_{u}=1.0\penalty 10000\ (m)$
+N avg. time (s) SR avg. time (s) SR
+9 32.71 0.50 54.91 1.00
+16 34.94 0.40 64.53 1.00
+Table II: The simulation results obtained in terms of time to reach the goal and success rate (SR) for $10$ instances in 22 different scenarios with $N=[9,16]$.
+We report the results with different parameter selection i.e., $d_{u}=0$ (m) and $d_{u}=1.0$ (m).
+For the comparison with PACNav [2], we simulate three scenarios with different number of robots and a varying number of randomly generated circular obstacles.
+In the simulation, we do not inject noise in the measurements, nevertheless tracking error is present.
 For each scenario, we perform $10$ runs for statistical evidence.
 The metrics that we took into account are the time required for all the robots to travel $42$ meters in the goal direction and the success rate in terms of mutual collision avoidance and obstacle avoidance.
-In Table II, we report the results obtained from PACNav [7] and from our approach.
-We considered the following scenario: $N=[4,9,9]$, $\delta_{i}=[0.25,0.5,0.5]$ (m), $\delta_{o}=[0.3,0.75,0.75]$ (m), and we selected the parameters as follows for all the three scenarios $r_{s,i}=4.5$ (m), $d_{1}=d_{3}=0.5$ (m), $d_{2}=d_{4}=1$ (m), $\beta_{i}^{D}=0.5$, $\Gamma_{ij}=8$ (m), $d_{u}=0.8$ (m).
-The adjacency matrix which defines the proximity maintenance constraints were selected as,
-$A_{d}=\left[\begin{smallmatrix}0&1&0&1\\ 1&0&1&0\\ 0&1&0&1\\ 1&0&1&0\end{smallmatrix}\right],\text{ and}\quad A_{d}=\left[\begin{smallmatrix}0&1&0&1&0&0&0&0&0\\ 1&0&1&0&1&0&0&0&0\\ 0&1&0&0&0&1&0&0&0\\ 1&0&0&0&1&0&1&0&0\\ 0&1&0&1&0&1&0&1&0\\ 0&0&1&0&1&0&0&0&1\\ 0&0&0&1&0&0&0&1&0\\ 0&0&0&0&1&0&1&0&1\\ 0&0&0&0&0&1&0&1&0\\ \end{smallmatrix}\right].$
-For our approach we considered two different values for the parameters $\epsilon=\epsilon_{p}=\epsilon_{o}$.
-The parameter selection for PACNav [7] has been done by trial and error, on the basis of the environment and the number of robots.
-Besides that, as seen in Table II, the approach does not scale in terms of performance (convergence, safety, and proximity maintenance) when the number of robots grows, due to the complex superimposition of forces given by the surroundings.
+In Table III, we report the results obtained from PACNav [2] and from our approach.
+We considered the following scenarios: $N=[4,9,9]$, $\delta_{i}=[0.25,0.5,0.5]$ (m), $\delta_{o}=[0.3,0.75,0.75]$ (m), and we select the parameters as follows for all the three scenarios $r_{s,i}=4.5$ (m), $d_{1}=d_{3}=0.5$ (m), $d_{2}=d_{4}=1$ (m), $\beta_{i}^{D}=0.5$, $\Gamma_{ij}=8$ (m), $d_{u}=0.8$ (m).
+The adjacency matrices defining the proximity maintenance constraints were selected as grid adjacency matrices.
+For our approach, we considered two different values for the parameters $\epsilon=\epsilon_{p}=\epsilon_{o}$.
+The parameter selection for PACNav [2] has been done by trial and error approach, on the basis of the environment and the number of robots.
+Besides that, as seen in Table III, PACNav does not scale in terms of performance (convergence, safety, and proximity maintenance) when the number of robots grows, due to the complex superimposition of forces given by the surroundings.
 PACNav [2] Our ($\epsilon=2.00$) Our ($\epsilon=1.052$)
@@
 3 9 - 0.00 120.33 1.00 55.87 1.00
-Table II: The simulation results obtained in terms of time to reach the goal and success rate for $10$ instances in 33 different scenarios, comparing PACNav [7] with our method in two different parameter configurations.
-For further simulation results, please refer to the attached multimedia material, where we include simulations with up to $16$ UAVs in crowded conditions.
-Notice that even if the sufficient condition in Proposition 1 is not respected, the robots can still safely reach the goal in environments with $56\%$ of free space, as shown in multimedia material.
+Table III: The simulation results obtained in terms of time to reach the goal and success rate for $10$ instances in 33 different scenarios, comparing PACNav [2] with our method in two different parameter configurations.
+For further simulation results, please refer to the multimedia material.
 V Experimental Results
@@
 The goal positions are set to be $40$ (m) ahead with respect to the initial position of each robot, notice that we do not need to have a common reference frame.
-However, we need to satisfy the requirement stated in Proposition 1, i.e., $\Gamma_{ij}>\|e_{i}-e_{j}\|>\Delta_{ij}$, $\forall i,j=1\dots N$ with $i\neq j$.
+We choose to adopt a fixed value of $d_{u}$ in the experiments.
+Due to the high measurements’ uncertainties the value of $d_{u}$ did not satisfy (15) all the time.
+In this regard, two practical solutions were adopted in the experiments: (i) We discard measurements with values of $\lambda^{\max}>15$, and (ii) if the measured distance between $i$-th and $j$-th robot with $j\in\bar{\mathcal{N}_{i}}$ is greater than $\Gamma_{ij}$, we re-project the position of the $j$-th robot to be at a distance less than $\Gamma_{ij}$ from the $i$-th robot, in order to ensure the existence of a nonempty set $\mathcal{F}_{i}$.
+These two strategies help to practically manage large estimation errors that exceed the theoretical threshold e.g., temporary failures in the estimation of distance to other robot’s due to tree occlusions.
 We tested the algorithm with up to $N=4$, by selecting $\delta_{i}=0.3$ (m), $\delta^{o}=0.05$ (m), $r_{s,i}=6$ (m), $d_{1}=d_{3}=0.5$ (m), $d_{2}=d_{4}=1$ (m), $\beta_{i}^{D}=2.5$, $\Gamma_{ij}=12$ (m), $d_{u}=0.8$ (m), $\epsilon_{p}=\epsilon_{o}=1.052$, and the adjacency matrix that defines the flocking constraints is $A_{d}=\left[\begin{smallmatrix}0&1&0&1\\ 1&0&1&0\\ 0&1&0&1\\ 1&0&1&0\end{smallmatrix}\right]$.
@@
 In Figure 7, we depict the distances between UAV $2$ and the other robots according to the UVDAR measurements.
-In Table III, we report the average velocities, the maximum velocities, and the time to reach the goal for each UAV during the mission.
+In Table IV, we report the average velocities, the maximum velocities, and the time to reach the goal for each UAV during the mission.
 To verify the success of the mission, we report in Figure 6-(b) the actual trajectories for all the 4 robots in the mission, while in Figure 8 we show the distances between the UAVs during the mission.
-As we expected the safety constraints and the flocking constraints (according to the adjancency matrix $A_{d}$) are satisfied.
-To show reliability and robustness to parameter changes, we report in Table IV the results obtained from other four experiments in the forest, with $N=3$ UAVs.
-Since the system is intrinsically chaotic it does not make sense to compare the results obtained, even if, as we can expect, a relation between $\beta^{D}$ and the time to reach the goal can be inspected from Table IV.
+The safety constraints and the flocking constraints (according to the adjancency matrix $A_{d}$) were satisfied during the entire mission.
+To show reliability and robustness to parameter changes, we report in Table V the results obtained from other four experiments in the forest, with $N=3$ UAVs.
+Since the system is intrinsically chaotic it does not make sense to compare the results obtained, even if, as we can expect, a relation between $\beta^{D}$ and the time to reach the goal can be inspected from Table V.
 Nevertheless, through the table we want to show the repeatability of the experiments and the robustness to parameter changes.
@@
 time (s) 150.61 174.95 159.20 174.20
-Table III: Average speed, maximum speed, and time to reach the goal of each UAV in the experiment from Figure 8.
-Figure 6: Distances measured by UVDAR system on UAV 2 from UAV 1 (blue), UAV 3 (orange), and UAV 4 (red). The flocking and safety constraints are in magenta.
-Figure 7: Distances between UAVs computed based on odometries transformed to a common reference frame obtained through alignment of maps of individual UAVs in a post-prossessing phase.
+Table IV: Average speed, maximum speed, and time to reach the goal of each UAV in the experiment from Figure 6.
 Refer to caption Refer to caption
@@
 The map was built by UAV 11.
+Figure 7: Distances measured by UVDAR system on UAV 2 from UAV 1 (blue), UAV 3 (orange), and UAV 4 (red). The flocking and safety constraints are in magenta.
+Figure 8: Distances between UAVs computed based on odometries transformed to a common reference frame obtained through alignment of maps of individual UAVs in a post-prossessing phase.
+The flocking and safety constraints are in magenta.
+According to $A_{d}$ the flocking constraints are not active for the pair of robots 22-44 and 11-33.
 UAV1 UAV2 UAV3
@@
 time (s) 82.49 80.33 62.30
-Table IV: Comparison of average speed, maximum speed, and time to reach the goal in four different experiments in the forest, showing repeatability and robustness to parameters changes.
+Table V: Comparison of average speed, maximum speed, and time to reach the goal in four different experiments in the forest, showing repeatability and robustness to parameters changes.
 VI Conclusions
-In this paper, we presented a novel distributed algorithm to swarm reliably and effectively in dense cluttered environments represented by a forest, accounting for tracking errors and measurements uncertainties.
-We provide sufficient conditions for the theoretical guarantees of safety, proximity maintenance and convergence, compare it with state of the art algorithms, and finally we validate our solution through experiments in the field and in dense environment such as a forest.
+In this paper, we presented a novel distributed algorithm to swarm reliably and effectively in dense cluttered environments represented by a forest.
+We provide sufficient conditions for safety and proximity maintenance.
+We propose an adaptive strategy to compensate for tracking errors and measurement uncertainties.
+We compare our approach with state-of-the-art algorithms, and finally we validate our solution through simulations and experiments in the field and in dense environment such as a forest.
 Our approach presents some limitations; the algorithm is limited to work on a plane, i.e., the reference in the altitude for each robot is constant and equal.
-To ensure the convergence towards the goal region, we assume that the individual obstacles are separated by a certain minimum distance (see Proposition 1, notice that these are just sufficient conditions and not necessary).
+The convergence towards the goal depends on both the environment of operation and the uncertainties taken into account.
 Providing to the algorithm a higher level re-planner can be beneficial to enhance performance.
-Future research directions will be focused on an effective 3D extension of the Lloyd algorithm and to increase the agility of the swarm.
+Future research directions will be focused on an effective 3D extension of the Lloyd algorithm and to further increase the agility of the swarm.
 CNN